### v1

# Multi-Head RAG: Solving Multi-Aspect Problems with LLMs

Abstract
Retrieval Augmented Generation (RAG) enhances the abilities of Large Language Models (LLMs) by enabling the retrieval of documents into the LLM context to provide more accurate and relevant responses. Existing RAG solutions do not focus on queries that may require fetching multiple documents with substantially different contents. Such queries occur frequently, but are challenging because the embeddings of these documents may be distant in the embedding space, making it hard to retrieve them all. This paper introduces Multi-Head RAG (MRAG), a novel scheme designed to address this gap with a simple yet powerful idea: leveraging activations of Transformer’s multi-head attention layer, instead of the decoder layer, as keys for fetching multi-aspect documents. The driving motivation is that different attention heads can learn to capture different data aspects. Harnessing the corresponding activations results in embeddings that represent various facets of data items and queries, improving the retrieval accuracy for complex queries. We provide an evaluation methodology and metrics, synthetic datasets, and real-world use cases to demonstrate MRAG’s effectiveness, showing improvements of up to 20% in relevance over standard RAG baselines. MRAG can be seamlessly integrated with existing RAG frameworks and benchmarking tools like RAGAS as well as different classes of data stores.
1 Introduction
Large Language Models (LLMs) transformed many machine learning tasks using in-context learning abilities. They achieved such accuracy by leveraging an increasing number of parameters, which in recent models have grown to hundreds of billions, making LLM training expensive in terms of both time and resources. It also comes with the danger of leaking confidential data into model weights [41, 34, 29]. Additionally, continuous training through fine-tuning is necessary to keep LLMs up-to-date. Even using the newest data, LLMs display an ongoing problem of hallucinations [45, 39, 14] by providing factually incorrect information. Retrieval Augmented Generation (RAG) was proposed [19, 12] in order to address these issues as well as others and make LLMs more trustworthy.
The key idea behind RAG is to enhance the generative model’s capabilities by integrating a retrieval system that can fetch relevant documents or passages from a large corpus of data. In this setting, when a query is received, the retrieval system first identifies and retrieves pertinent information, which is fed into the generative model’s context for a more accurate and relevant response. Instead of the model storing information within its weights, RAG effectively leverages external knowledge, reducing hallucinations (by grounding the LLM reply in reliable sources), and ensuring that responses contain up-to-date knowledge (e.g., by accessing the Internet), all without requiring expensive training.
More specifically, there are two main stages in a RAG pipeline: data preparation and query execution. During data preparation, one constructs a vector database (DB) populated with embeddings and their corresponding data items such as documents. During query execution, one constructs an embedding of that query and retrieves data items in the store with similar embeddings.
Intense recent research efforts have been put into RAG [11, 46, 13, 15, 42, 26, 21]. On one hand, different RAG designs have been proposed, for example RAPTOR [32], Self-RAG [3], Chain-of-Note [43], and many others [2, 7, 8, 24, 44, 36, 40]. In general, these schemes focus on making the retrieved data more accurate and relevant to the query. On the other hand, there have also been efforts into benchmarking and datasets for RAG evaluation [5, 37, 22, 9].
Despite all these advances, we observe that no existing RAG scheme or evaluation methodology explicitly targets an important class of problems that come with a high degree of multi-aspectuality. These are problems that require combining several (potentially many) significantly different aspects in a single query. As a simple illustrative example of such a query, consider the question “What car did Alexander the Great drive?”, and assume that the queried model has not been trained on history. When using RAG, to answer this question accurately, one would retrieve two documents, one describing Alexander the Great and one outlining the history of car manufacturing. However, the embeddings of these two documents could be far away from each other in the embedding space. At the same time, such queries are common in different industry settings, as indicated by extensive discussions with our industry collaborators. Imagine a chemical processing plant experiencing an equipment accident. One could use an LLM to find the accident cause, which might require the retrieval of multiple, potentially confidential documents to provide the necessary context. These documents could be related to different aspects, for example psychological profiles of workers (“Was the accident due to mismanaging a worker?”), equipment purchase records (“Was some equipment part too old?”), maintenance (“Was some equipment part rusty?”), weather (“Was there a particularly strong thunderstorm at the accident time that could have caused dangerous power spikes in the grid?”), or even microclimate (“Was it too humid for an extended period of time in the production hall?”). As we illustrate in Section 4, such problems pose challenges for existing RAG schemes and have been unaddressed by modern RAG benchmarking pipelines.
Figure 1: An overview of the decoder architecture, and a comparison of how standard RAG and Multi-Head RAG embeddings are generated.
In this work, we propose Multi-Head RAG (MRAG): a scheme that addresses the above problem. Common practice in modern RAG designs is the use of embeddings based on last-layer decoder block activations. Our key idea is to use instead the activations of the multi-head attention part of the decoder block as embeddings. The Transformer architecture can be seen as a pipeline with many (e.g., 96 for GPT-3 [6]) blocks, where a single block consists of an attention module and a feed-forward module. Each individual attention module is multi-headed: it consists of multiple parts called heads that learn different sets of weight matrices; see Figure 1 for an overview. It is conjectured that these different heads could capture different aspects of the processed data. We use this as a driving design feature that facilitates capturing the potential multi-aspectuality of the data without increasing space requirements compared to standard RAG (contribution 1).
Such multi-aspect embeddings are then directly used for both data items and query representation. Considering multi-aspectuality explicitly comes with challenges. For example, how to assess the effectiveness of a RAG solution in retrieving data that indeed does cover multiple aspects of a given domain. For this, we establish an evaluation methodology as well as a full data construction and query processing pipeline that implements the multi-aspect embedding idea (contribution 2). Our datasets facilitate broad evaluation by considering both fully-automatically generated, synthetic data and analyzing specific industry use cases that show the benefits of MRAG (contribution 3). Our evaluation illustrates the benefits in the relevance of retrieved documents, for example 20% over a modern RAG baseline for fetching multi-aspect Wikipedia articles (contribution 4). We also show how MRAG and its benchmarking principles can be seamlessly integrated with both existing RAG solutions and benchmarking frameworks such as RAGAS (contribution 5). MRAG’s code is publicly available [^1] https://github.com/spcl/MRAG.
2 The MRAG Formulation & Pipeline
We now present in detail the mathematical underpinning of MRAG and its corresponding pipeline.
[…]
Multi-Head RAG Formulation The key idea behind MRAG is simple: instead of the single activation vector generated by the last feed-forward decoder layer for the last token, we harness the $H$ separate activation vectors generated by the last attention layer for the last token, before processing it via $\mathbf{W_{o}}$. This can be formulated as a set of embeddings $\mathcal{S}=\{\mathbf{e}_{k}\forall_{k}\}$ where $\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})$, which is simply the set of all outputs from the attention heads on the last token $\mathbf{x}_{n}$ of the input. As processing with multiple heads does not change the size of the output vector, $\mathcal{S}$ has the same space requirements as standard RAG. However, because we capture the separate embeddings before their mixing with $\mathbf{W}_{o}$, we conjecture that it gives more information about what the different parts of the input attend to, facilitating capturing multi-aspectuality.
Naming We use the terms “single-aspect embedding” and “multi-aspect embedding” to refer to, respectively, a small embedding extracted from a single attention head and a collection of all single-aspect embeddings extracted from an attention layer.
2.1 Overview of the Multi-Head RAG Pipeline
We now describe how the above embedding model fits the RAG pipeline. Figure 2 shows a summary of the design. The MRAG pipeline consists of two main parts, dedicated to data preparation [Uncaptioned image] and query execution [Uncaptioned image]. Both parts heavily use the data store [Uncaptioned image] (vector DB).
2.1.1 Data Preparation
During data preparation [Uncaptioned image], we populate a data store [Uncaptioned image] with multi-aspect MRAG text embeddings [Uncaptioned image] and their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image] (MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data). We create the multi-aspect embedding [Uncaptioned image] of each text chunk [Uncaptioned image] using a selected decoder-based embedding model [Uncaptioned image] (this part is detailed in Section 2.2). The user of the pipeline can plug in their model [Uncaptioned image] of choice as well as use their input data. We also offer a dedicated synthetic data generator [Uncaptioned image] that can be used to construct multi-aspect input documents [Uncaptioned image] (we detail this part in Section 3) for evaluation purposes.
MRAG stores data differently than standard RAG, where a single embedding [Uncaptioned image] points to a single text chunk [Uncaptioned image]. For MRAG, each multi-aspect embedding consists of $h$ single-aspect embeddings [Uncaptioned image], each pointing to the original text chunk [Uncaptioned image]. So the data store [Uncaptioned image] contains $h$ embedding spaces, each capturing a different aspect of the text. This crucial feature allows MRAG to compare query [Uncaptioned image] and text chunks [Uncaptioned image] in multiple embedding spaces that capture multiple aspects of the data.
2.1.2 Query Execution
During query execution [Uncaptioned image], we first generate a multi-aspect embedding [Uncaptioned image] of the input query [Uncaptioned image], using the selected embedding model [Uncaptioned image] (details in Section 2.2). Then, we find the nearest multi-aspect embeddings [Uncaptioned image] and their corresponding text chunks [Uncaptioned image] in the data store [Uncaptioned image] using a special multi-aspect retrieval strategy [Uncaptioned image] (detailed in Section 2.3). Finally, the retrieved data can optionally be assessed [Uncaptioned image] with novel metrics regarding how well it corresponds to the multi-aspect requirements (detailed in Section 3). As with the data preparation [Uncaptioned image] stage, the query execution [Uncaptioned image] stage is flexible, and the user can plug in their models [Uncaptioned image] / [Uncaptioned image] of choice and use their own queries [Uncaptioned image]. We also offer a dedicated synthetic query generator [Uncaptioned image] that can be used to construct multi-aspect input queries [Uncaptioned image] (detailed in Section 3) for evaluation purposes.
Figure 2: Overview of the MRAG pipeline, consisting of two parts: data preparation Refer to caption and query execution Refer to caption. The embedding model Refer to caption and the data store Refer to caption are used by both parts. The data store Refer to caption contains text embeddings Refer to caption linking to text chunks Refer to caption reflecting three different aspects (cyan, magenta, yellow). Blocks marked by a star Refer to caption are a novelty of this work.
2.2 Constructing Multi-Aspect Embeddings [Uncaptioned image]
MRAG can leverage any embedding model with multi-head attention support to construct the multi-aspect embeddings for a given input text. In this work, we consider two embedding models from the MTEB leaderboard [16] as potential candidates. Specifically, the SFR-Embedding-Model [25] and the e5-mistral-7b-instruct [35], both based on the Mistral 7B architecture with 32 decoder blocks and 32 attention heads per multi-head attention.
While our approach allows for extracting and using the multi-aspect embeddings from any decoder block, and from different layers within a block, we found that multi-aspect embeddings extracted from the last multi-head attention worked best in our experimental setting. We provide further discussion on the carried out experiments in Section 4.
2.3 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
A retrieval strategy determines how we select the closest text chunks from the DB given a multi-aspect embedding of the user query. In general, the MRAG retrieval strategy consists of three steps. First, during data preparation, we assign importance scores to all $h$ embedding spaces. Intuitively, these scores capture the fact that different spaces (and the corresponding heads) may be more or less relevant for the used data. Then, during query execution, MRAG starts by applying the traditional RAG retrieval separately for each embedding space. This returns a list of $c$ closest text chunks for each embedding space (a total of $h$ lists). Here, we use a special voting strategy to pick overall top $k$ out of all $hc$ chunks, using the pre-computed importance scores.
Algorithm 1 Importance scores for heads.
  for each head $h_{i}$ do
     $a_{i}\leftarrow 0$; $b_{i}\leftarrow 0$
     $count\_a_{i}\leftarrow 0$; $count\_b_{i}\leftarrow 0$
     for each embedding $e_{ij}$ in $h_{i}$ do
        $a_{i}\leftarrow a_{i}+||e_{ij}||$
        $count\_a_{i}\leftarrow count\_a_{i}+1$
        for each embedding $e_{ih}$ do
           $b_{i}\leftarrow b_{i}+\text{cosine-distance}(e_{ij},e_{ih})$
           $count\_b_{i}\leftarrow count\_b_{i}+1$
        end for
     end for
     $a_{i}\leftarrow a_{i}/count\_a_{i}$; $b_{i}\leftarrow b_{i}/count\_b_{i}$
     $s_{i}\leftarrow a_{i}\cdot b_{i}$
  end for
Algorithm 1 details the construction of importance scores. It is a heuristic based on extensive empirical evaluation; it gives high-quality results across the tested datasets and tasks. Intuitively, the score $s_{i}$ of a given head $h_{i}$ consists of two parts, $a_{i}$ and $b_{i}$. $a_{i}$ is the average of L2 norms of all embeddings in the vector space $i$; it represents how important a given head is: the larger the norms, the more attention was given to this attention head. $b_{i}$ is the average of cosine distances between all (or a randomly sampled subset, if the user wants to reduce pre-compute time) embeddings in vector space $i$. Intuitively, $b_{i}$ is a proxy for measuring the “spread” of vector space $i$: the larger $b_{i}$, the larger the average angle between different embeddings in this space is. Deriving $s_{i}$ as a product $a_{i}\cdot b_{i}$ ensures that we reward heads with high average attention and high average spread, but simultaneously penalize heads with lower average attention or with low average spread (both $a_{i}$ and $b_{i}$ are appropriately scaled).
The used voting strategy combines the constructed lists of text chunks from individual embedding spaces into a single list of top $k$ chunks. The strategy is very simple (the corresponding Algorithm 2 is in the Appendix). Each text chunk from a list $i$ of the vector space $i$ has a certain position on this list, we denote this position with $p$. We obtain a weight for this chunk as $s_{i}\cdot 2^{-p}$; $s_{i}$ is the previously defined importance score of the space $i$. Multiplying $s_{i}$ with $2^{-p}$ exponentially lowers the significance of less relevant text chunks. Finally, all chunks from all lists are sorted using their weights and the top $k$ chunks form the final list.
2.3.1 Integration with Data Stores
MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches. It can be combined with both the exact and the approximate NN to find the matching (embedding, chunk)-pairs. These two parts of the broader RAG processing pipeline are orthogonal to MRAG.
3 Multi-Aspect Datasets, Queries, and Metrics
To assess how well MRAG performs on multi-aspect queries, and to compare it to modern RAG schemes, we need (1) datasets of documents that capture multi-aspectuality, (2) queries to the LLM that touch upon multi-aspectuality and require retrieving documents from the multi-aspect dataset, and (3) metrics that assess how well a RAG scheme retrieves such multi-aspect data. We now describe these three elements. In Section 4, we also briefly discuss real-world data and queries used.
Multi-Aspect Dataset Generation We first select conceptually different categories of documents. We primarily focus on publicly available Wikipedia articles and select 25 categories (e.g., countries, board games, historical swords, shipwrecks, etc.). For each category, we sample 50 documents. The first part of the document (overview) is used as a text chunk to be embedded. We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes.
Multi-Aspect Query Generation We also require queries that touch upon a given number of $n$ aspects. For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories. We create such queries by selecting $n$ categories, sampling a document from each selected category (ensuring there are no duplicates overall), and then generating a story that combines these documents, using an LLM (GPT-3.5 Turbo). We construct 25 queries with 1, 5, 10, 15 and 20 aspects (125 queries in total). An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 3.
Metrics We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality. For a query $Q$, a used retrieval strategy $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$. Then, $S(Q,n)$ is the set of the actually retrieved documents. We define the Retrieval Success Ratio as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, i.e., the ratio of successfully retrieved relevant documents. Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category. To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$. It has the same form as $\Xi(Q,n)$ above, with one difference: $S(Q,n)$ is now the set of all the retrieved documents that belong to categories of the ideal desired documents. Finally, to combine these two metrics, we use the Weighted Retrieval Success Ratio $\Xi_{w}$ as $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$. By varying $w$, the user can adjust the importance of exact document matches and category matches. An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 3.
4 Evaluation
We now illustrate the advantages of MRAG over the state of the art.
Comparison Baselines We compare MRAG to two main baselines: Standard RAG and Split RAG. The first represents a modern RAG pipeline in which each document uses the activations of the last decoder layer as its embedding. The second is a blend between Standard RAG and MRAG. Specifically, it splits the activation of the last decoder layer in the same way as MRAG and applies a voting strategy. The purpose of Split RAG is to show that MRAG’s benefits come from using the multi-head output as embedding and not merely using multiple embedding spaces. Additionally, we consider Fusion RAG [30], an optional mechanism that we harness to further enhance the benefits of MRAG at the cost of additional tokens (detailed in Section 4.2).
We use queries and metrics introduced in Section 3. We use the weighted retrieval success ratio with 2:1 weighting, which considers category matches as relevant but prioritizes the exact document matches. Figure 3 shows an example query and metrics usage. Each query requires retrieving a specific number of documents and the corresponding non-overlapping categories which define the ground truth. We fetch the top $k$ documents from a database, where $k$ is the “total number of documents fetched for a tested RAG scheme” (including potentially mismatches). Among these $k$ documents, we search for matches with the ground truth.
Figure 3: An example query used to evaluate different RAG strategies. We mention the documents to be fetched in the text and then assess the success ratio of different RAG strategies in finding these documents and their categories. We mark exact document matches Refer to caption , category matches Refer to caption , documents that match a category multiple times Refer to caption , and text segments with no matching document Refer to caption . Finally, we show the weighted success ratio for each strategy, taking a 2:1 weighting (prioritizing the exact article matches).
Samples & Summaries Each data point in our plots corresponds to 25 queries. We present the data using standard boxplots to showcase the distribution. Our primary focus is on the average retrieval performance among those 25 queries.
4.1 Analysis of Results
We start from the query example in Figure 3 and show first the absolute retrieval performance of MRAG over Standard RAG in Figure 4. We fix the number of aspects present in the queries to 10, and vary the total number of retrieved documents from 10 to 30. MRAG consistently outperforms Standard RAG ($>10\%$ increase in the retrieval success ratio on average for exact document matches). Moreover, the retrieval performance increase is even more significant on category matches ($>25\%$ increase in the retrieval success ratio on average). The performance increase is further detailed in the histograms on the right side. Here, for a specific number of documents fetched, MRAG’s histogram indicates a better distribution of retrieval success ratios (across all 25 queries).
Next, Figure 5 shows the relative weighted performance improvement of MRAG with respect to Standard RAG as we vary the number of aspects present in the queries. We show data for two different embedding models (SFR and e5). MRAG consistently outperforms the Standard RAG by 10-20% on average, not only across the number of documents fetched, but also across the number of aspects present in the replies, for both models.
Figure 4: Retrieval success ratio over 25 queries between MRAG and Standard RAG, where each query includes 10 different aspects. The upper part presents exact document matches while the lower part presents category only matches (we explain the metrics used in Section 3). A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (the number of documents fetched for each query is 30).
Figure 5: Relative retrieval improvement of MRAG over Standard RAG across queries with different numbers of aspects and different embedding models (SFR in the left side, e5 in the right side).
[…]
Documents Fetched Multi-Aspect Dataset Legal Dataset Accidents Dataset
SFR e5 SFR SFR
MRAG Standard RAG | MRAG Standard RAG | MRAG Standard RAG | MRAG Standard RAG
1 24/25 25/25 24/25 25/25 24/25 24/25 25/25 25/25
2 25/25 25/25 25/25 25/25 25/25 25/25 25/25 25/25
3 25/25 25/25 25/25 25/25 25/25 25/25 25/25 25/25
Table 1: Retrieval success ratio (the exact document match) for 25 queries with a single aspect.
We additionally show in Table 1 that MRAG performs on-par with Standard RAG on queries from our multi-aspect dataset where only a single aspect is expected. Hence, our approach does not suffer from significant decrease in performance for single-aspect tasks.
4.2 Further Improvements with Additional Tokens
We now show that MRAG can be seamlessly integrated with other RAG approaches: We combine MRAG with Fusion RAG, representing RAG schemes that use an LLM (additional token cost) for more accurate retrieval. Fusion RAG uses an LLM to create a fixed number of questions about the RAG query. Each question is separately applied through an embedding model using Standard RAG. We apply MRAG’s approach to each of these questions and denote the combined scheme as Fusion MRAG. Red plots of Figure 6 show that both Fusion RAG and Fusion MRAG perform better than Standard RAG, on average gaining 10 to 30% in accuracy. Fusion MRAG performs consistently better than pure Fusion RAG, indicating that these optimizations can be combined together. However, both Fusion strategies introduce a greater variance than MRAG and additional costs in terms of compute, latency, and tokens.
4.3 Benefits from Multi-Head Attention Solely
We also compare MRAG to the Split RAG baseline in Figure 6. The blue plots show the relative weighted performance of MRAG and Split RAG over Standard RAG. MRAG performs better than Split RAG, illustrating that its high accuracy is due to the actual multi-head part, and not merely just partitioning the vector and using multiple embedding spaces.
Figure 6: Relative retrieval improvements of MRAG over Standard RAG for the SFR embedding model compared with Split RAG (the blue plots), and the relative retrieval improvements of Fusion MRAG over both Fusion RAG and MRAG (the red plots).
4.4 Real-World Workloads
Figure 7: Average improvement of the retrieval success ratio of MRAG and Split RAG over Standard RAG for two real-world workloads constructing legal documents (left) and discovering causes of industry accidents (right).
To further illustrate advantages of MRAG, we also consider two real-word use cases from in-house industry data analytics projects, namely, the synthesis of legal documents and the analysis of causes of chemical plant accidents. The results are in Figure 7. In the former (the left side), the task is to create a document based on user requirements that may be related to different aspects, for example to the law being considered (e.g., the British or the US one), the subject (e.g., energetic or civil), the style of the document (e.g., aggressive or mild), etc.. This task is executed with RAG that can fetch documents from a database. In the latter (the right side), the task is to discover a cause of an accident. Here, one also wants to retrieve documents from a database that should be used in the LLM context to facilitate discovering the cause of the accident. The causes are grouped in categories such as utility impact due to severe weather, lack of preparedness and planning, incorrect installation of equipment, lack of maintenance, etc.. Similarly to the previous analyses, we measure the retrieval success ratio over corresponding databases. MRAG offers advantages over other schemes.
4.5 Additional Analyses
Figure 8: Evaluation of different voting strategies for MRAG and Split RAG
We also analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one) and the impact of using different voting strategies for MRAG as well as for Split RAG.
We consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention. We discover that the last multi-head attention performs the best when compared with the Standard RAG.
We also illustrate selected representative data from a long investigation two additional voting strategies for MRAG. We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$. Figure 8 shows that these voting strategies perform worse on average than our selected strategy for MRAG, justifying its design and selection (described in Section 2.3).
We also consider two voting strategies for Split RAG, to further deepen the empirical evaluation. Split (1) only uses the exponential lowering of significance ($w_{i,p}=2^{-p}$) and Split (2) which uses the same strategy as MRAG ($w_{i,p}=s_{i}\cdot 2^{-p}$). Figure 8 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
5 Related Work
Our work touches on many areas which we now briefly discuss.
Many RAG schemes appeared recently [11], using the output of the last decoder layer for embedding generation. In contrast, MRAG leverages different embedding spaces of attention heads to focus on different aspects of documents and queries. As such, it can be combined with other schemes to further improve RAG pipelines.
Retrieval is sometimes enhanced by a cross-encoder reranking phase [31, 27, 28, 20, 10, 23]. In such solutions, typically after retrieving a set of relevant chunks, they are re-ranked using specialized models. In this work, we focus solely on the first retrieval phase, so MRAG can be seamlessly used in conjunction with such cross-encoders.
Structure-enhanced RAG schemes employ different strategies for structuring text to improve retrieval quality. A common idea is to construct a Knowledge Graph from text, which enables retrieval amongst entities and relationships [18, 7, 17, 4, 38]. RAPTOR [32] generates multi-level summaries for clusters of related chunks, building a tree of summaries with increasing levels of abstraction to better capture the meaning of the text. Graph RAG [8] creates a Knowledge Graph, and summarizes communities in the graph, which provide data at the different levels of abstraction. All these systems try to improve RAG quality by utilizing additional structures that describe entity relationships or the inner organization of text. Usually, they need a sophisticated preprocessing phase to prepare such structures. MRAG achieves the improvement solely based on the embedding model and has no additional storage requirements, and can be combined with any of these schemes.
6 Conclusion
Retrieval Augmented Generation (RAG) is pivotal for democratizing access to accurate and relevant outputs from large language models (LLMs). Enhancing the precision and relevance of these outputs is a critical goal, especially given the challenges posed by queries requiring the retrieval of multiple documents with significantly different contents. These complex queries are common across various domains, but existing RAG solutions struggle because the embeddings of the necessary documents can be far apart in the embedding space, complicating their retrieval.
To address this gap, we introduced Multi-Head RAG (MRAG), a novel scheme that leverages the activations from the multi-head attention layer of decoder models instead of the traditional feed-forward layer. This approach is grounded in the insight that different attention heads can capture distinct aspects of the data. By using these diverse activations, MRAG creates embeddings that better represent the multifaceted nature of data items and queries, thus enhancing the retrieval accuracy for complex, multi-aspect queries. The simplicity and versatility of this idea allow it to be seamlessly integrated into any modern RAG pipeline or data analytics framework.
Our comprehensive evaluation methodology, including specific metrics, synthetic datasets, and real-world use cases, demonstrates MRAG’s effectiveness. The results indicate a significant improvement in the relevance of retrieved documents, with up to 20% better performance compared to modern RAG baselines. This validates MRAG’s potential to handle the intricacies of multi-aspect queries effectively.
Moreover, MRAG proves to be both cost-effective and energy-efficient. It does not require additional LLM queries, multiple model instances, increased storage, or multiple inference passes over the embedding model. This efficiency, combined with the enhanced retrieval accuracy, positions MRAG as a valuable advancement in the field of LLMs and RAG systems. By addressing the challenges of multi-aspectuality in queries, MRAG paves the way for more reliable and accurate LLM applications across diverse industries.
Acknowledgments and Disclosure of Funding
We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines, and for their excellent technical support. We thank Timo Schneider for help with infrastructure at SPCL. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project was supported by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei Technologies. This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017103.
Appendix
Appendix A Model Design: Additional Details
A.1 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
Algorithm 2 Voting strategy.
  $l\leftarrow[]$
  for each head $h_{i}$ and its score $s_{i}$ do
     find best matching $k$ text chunks
     for each chunk $d_{i,p}$ with index $p$ in top $k$ do
        $w_{i,p}\leftarrow s_{i}\cdot 2^{-p}$
        add tuple $(d_{i,p},w_{i,p})$ to $l$
     end for
  end for
  sort $l$ using weights $w_{i,p}$; return top $k$ elems
Appendix B Evaluation Methodology: Additional Details
B.1 Compute Resources
Our experiments were executed with compute nodes containing 4x NVIDIA GH200 and a total memory of 800 GB. In general one GPU with at least 40GB of memory should suffice. We used at most 50GB of storage and the OpenAI API as an external resource. The full experiments took at most three hours of GPU time and the cost for the OpenAI API were at most \$15. We carried out additional experiments, which amounted to around 20 hours of GPU time and cost of \$25 for the OpenAI API. Additional evaluation was executed with a mix of compute resources including NVIDIA A100 and V100 GPUs.
B.2 Dataset Details
Table 2: Prompt template for query generation.
Please create a story about the attached <number of articles> articles on the topics <list of titles>.
It is very important that each of the attached articles is relevant to the story, in a way that references the content of the article, not just its title. But please also mention each title at least once. Please make sure that all of the attached articles are relevant to your story, and that each article is referenced in at least two sentences! They do not necessarily have to be referenced in the same order, but make sure no article is forgotten.
Important: Output only the story, no additional text. And do not use bullet points, or paragraphs.
Articles:
Article <title>:
<body>
Again, make sure that you reference all the following topics in your story: <list of titles>

### v2

@@
 Harnessing the corresponding activations results in embeddings that represent various facets of data items and queries, improving the retrieval accuracy for complex queries.
-We provide an evaluation methodology and metrics, synthetic datasets, and real-world use cases to demonstrate MRAG’s effectiveness, showing improvements of up to 20% in relevance over standard RAG baselines.
+We provide an evaluation methodology and metrics, multi-aspect datasets that we release online, and real-world use cases to demonstrate MRAG’s effectiveness, showing improvements of up to 20% in relevance over standard RAG baselines.
 MRAG can be seamlessly integrated with existing RAG frameworks and benchmarking tools like RAGAS as well as different classes of data stores.
@@
 They achieved such accuracy by leveraging an increasing number of parameters, which in recent models have grown to hundreds of billions, making LLM training expensive in terms of both time and resources.
-It also comes with the danger of leaking confidential data into model weights [41, 34, 29].
+It also comes with the danger of leaking confidential data into model weights (Yan et al., 2024; Shi et al., 2024; Patil et al., 2024).
 Additionally, continuous training through fine-tuning is necessary to keep LLMs up-to-date.
-Even using the newest data, LLMs display an ongoing problem of hallucinations [45, 39, 14] by providing factually incorrect information.
-Retrieval Augmented Generation (RAG) was proposed [19, 12] in order to address these issues as well as others and make LLMs more trustworthy.
-The key idea behind RAG is to enhance the generative model’s capabilities by integrating a retrieval system that can fetch relevant documents or passages from a large corpus of data.
+Even using the newest data, LLMs display an ongoing problem of hallucinations (Zhang et al., 2023; Xu et al., 2024c; Huang et al., 2023) by providing factually incorrect information.
+Retrieval Augmented Generation (RAG) was proposed (Lewis et al., 2020; Guu et al., 2020) in order to address these issues as well as others and make LLMs more trustworthy.
+The key idea behind RAG is to enhance the generative model’s capabilities by integrating a retrieval system that fetches relevant passages from a large corpus of data.
 In this setting, when a query is received, the retrieval system first identifies and retrieves pertinent information, which is fed into the generative model’s context for a more accurate and relevant response.
@@
 During query execution, one constructs an embedding of that query and retrieves data items in the store with similar embeddings.
-Intense recent research efforts have been put into RAG [11, 46, 13, 15, 42, 26, 21].
-On one hand, different RAG designs have been proposed, for example RAPTOR [32], Self-RAG [3], Chain-of-Note [43], and many others [2, 7, 8, 24, 44, 36, 40].
+Intense recent research efforts have been put into RAG (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2024; Huang & Huang, 2024; Yu et al., 2024; Mialon et al., 2023; Li et al., 2022).
+On one hand, different RAG designs have been proposed, for example RAPTOR (Sarthi et al., 2024), Self-RAG (Asai et al., 2023), Chain-of-Note (Yu et al., 2023), and many others (Abdallah & Jatowt, 2024; Delile et al., 2024; Edge et al., 2024; Manathunga & Illangasekara, 2023; Zeng et al., 2024; Wewer et al., 2021; Xu et al., 2024b).
 In general, these schemes focus on making the retrieved data more accurate and relevant to the query.
-On the other hand, there have also been efforts into benchmarking and datasets for RAG evaluation [5, 37, 22, 9].
+There have also been efforts into benchmarking and datasets for RAG evaluation (Chen et al., 2024b; Xiong et al., 2024; Lyu et al., 2024; Es et al., 2023).
 Despite all these advances, we observe that no existing RAG scheme or evaluation methodology explicitly targets an important class of problems that come with a high degree of multi-aspectuality.
@@
 Our key idea is to use instead the activations of the multi-head attention part of the decoder block as embeddings.
-The Transformer architecture can be seen as a pipeline with many (e.g., 96 for GPT-3 [6]) blocks, where a single block consists of an attention module and a feed-forward module.
+The Transformer architecture can be seen as a pipeline with many (e.g., 96 for GPT-3 (Wang et al., 2024b)) blocks, where a single block consists of an attention module and a feed-forward module.
 Each individual attention module is multi-headed: it consists of multiple parts called heads that learn different sets of weight matrices; see Figure 1 for an overview.
 It is conjectured that these different heads could capture different aspects of the processed data.
-We use this as a driving design feature that facilitates capturing the potential multi-aspectuality of the data without increasing space requirements compared to standard RAG (contribution 1).
+We use this as a driving design feature that facilitates capturing the potential multi-aspectuality of the data without increasing space requirements compared to standard RAG, and without any fine-tuning or other modifications to the harnessed model (contribution 1).
 Such multi-aspect embeddings are then directly used for both data items and query representation.
 Considering multi-aspectuality explicitly comes with challenges.
-For example, how to assess the effectiveness of a RAG solution in retrieving data that indeed does cover multiple aspects of a given domain.
+For example, it is unclear how to assess whether a RAG solution does indeed harness multiple aspects when fetching documents.
 For this, we establish an evaluation methodology as well as a full data construction and query processing pipeline that implements the multi-aspect embedding idea (contribution 2).
-Our datasets facilitate broad evaluation by considering both fully-automatically generated, synthetic data and analyzing specific industry use cases that show the benefits of MRAG (contribution 3).
-Our evaluation illustrates the benefits in the relevance of retrieved documents, for example 20% over a modern RAG baseline for fetching multi-aspect Wikipedia articles (contribution 4).
+Our datasets facilitate broad evaluation by considering both fully-automatically generated, synthetic data as well as data based on specific industry use cases that show the benefits of MRAG (contribution 3).
+We ensure the relevance of our RAG datasets in real use cases by working directly with tech leaders (e.g., a generative AI division head) from 3 corporations, all of which actively use RAG in their own LLM infrastructures.
+Our evaluation illustrates the benefits in the relevance of retrieved documents, for example 20% over a modern RAG baseline for fetching multi-aspect Wikipedia articles, and comparable performance for single-aspect queries (contribution 4).
 We also show how MRAG and its benchmarking principles can be seamlessly integrated with both existing RAG solutions and benchmarking frameworks such as RAGAS (contribution 5).
@@
 We now present in detail the mathematical underpinning of MRAG and its corresponding pipeline.
+Decoder Formulation We first introduce formally the decoder architecture.
+We omit, for clarity, unnecessary details such as layer normalizations.
+The input is a text chunk that consists of $n$ tokens.
+The output of attention head $h$ for the $i$th token $x_{i}$ is defined as follows (Vaswani et al., 2017):
+$\text{head}^{h}(\mathbf{x}_{i})=\sum_{j}w_{ij}\mathbf{v}_{j}^{h},\quad\text{where}$ (1)
+$\vspace{-0.5em}w_{ij}=\text{softmax}\left(\left(\mathbf{q}_{i}^{h}\right)^{T}\mathbf{k}_{j}^{h}\right),\quad\mathbf{q}_{i}^{h}=\mathbf{W}_{q}^{h}\mathbf{x}_{i},\quad\mathbf{k}_{j}^{h}=\mathbf{W}_{k}^{h}\mathbf{x}_{j},\quad\mathbf{v}_{j}^{h}=\mathbf{W}_{v}^{h}\mathbf{x}_{j}$ (2)
+where $\mathbf{W}_{q}^{h},\mathbf{W}_{k}^{h},\mathbf{W}_{v}^{h}$ are, respectively, learnable query, key, and value projections associated with head $h$, and $\mathbf{x}_{j}$ is the vector embedding of the $j$th token $x_{j}$.
+These outputs get combined to form the output of the $i$th multi-head attention block as follows:
 […]
+where matrix $\mathbf{W}_{o}$ is the linear layer that combines the outcomes of all the attention heads. This step is then followed by the Transformer feed-forward layer.
 […]
 Multi-Head RAG Formulation The key idea behind MRAG is simple: instead of the single activation vector generated by the last feed-forward decoder layer for the last token, we harness the $H$ separate activation vectors generated by the last attention layer for the last token, before processing it via $\mathbf{W_{o}}$.
-This can be formulated as a set of embeddings $\mathcal{S}=\{\mathbf{e}_{k}\forall_{k}\}$ where $\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})$, which is simply the set of all outputs from the attention heads on the last token $\mathbf{x}_{n}$ of the input.
+This can be formulated as a set of embeddings $\mathcal{S}$:
+$\mathcal{S}=\{\mathbf{e}_{k}\forall_{k}\}\quad\text{where}\quad\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})$ (4)
+which is simply the set of all outputs from the attention heads on the last token $\mathbf{x}_{n}$ of the input.
 As processing with multiple heads does not change the size of the output vector, $\mathcal{S}$ has the same space requirements as standard RAG.
@@
 2.1.1 Data Preparation
-During data preparation [Uncaptioned image], we populate a data store [Uncaptioned image] with multi-aspect MRAG text embeddings [Uncaptioned image] and their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image] (MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data).
+When preparing data [Uncaptioned image], we populate a data store [Uncaptioned image] with multi-aspect MRAG text embeddings [Uncaptioned image] and their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image] (MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data).
 We create the multi-aspect embedding [Uncaptioned image] of each text chunk [Uncaptioned image] using a selected decoder-based embedding model [Uncaptioned image] (this part is detailed in Section 2.2).
@@
 Then, we find the nearest multi-aspect embeddings [Uncaptioned image] and their corresponding text chunks [Uncaptioned image] in the data store [Uncaptioned image] using a special multi-aspect retrieval strategy [Uncaptioned image] (detailed in Section 2.3).
+We ensure that there is no overhead in latency due to multiple aspects because computing these different smaller embeddings is done fully in parallel.
 Finally, the retrieved data can optionally be assessed [Uncaptioned image] with novel metrics regarding how well it corresponds to the multi-aspect requirements (detailed in Section 3).
@@
 MRAG can leverage any embedding model with multi-head attention support to construct the multi-aspect embeddings for a given input text.
-In this work, we consider two embedding models from the MTEB leaderboard [16] as potential candidates.
-Specifically, the SFR-Embedding-Model [25] and the e5-mistral-7b-instruct [35], both based on the Mistral 7B architecture with 32 decoder blocks and 32 attention heads per multi-head attention.
+In this work, we consider two embedding models from the MTEB leaderboard (Huggingface, 2024) as potential candidates.
+Specifically, the SFR-Embedding-Model (Meng et al., 2024) and the e5-mistral-7b-instruct (Wang et al., 2024a), both based on the Mistral 7B architecture with 32 decoder blocks and 32 attention heads per multi-head attention.
 While our approach allows for extracting and using the multi-aspect embeddings from any decoder block, and from different layers within a block, we found that multi-aspect embeddings extracted from the last multi-head attention worked best in our experimental setting.
@@
 Here, we use a special voting strategy to pick overall top $k$ out of all $hc$ chunks, using the pre-computed importance scores.
+Algorithm 1 details the construction of importance scores.
+It is a heuristic based on extensive empirical evaluation; it gives high-quality results across the tested datasets and tasks.
+Intuitively, the score $s_{i}$ of a given head $h_{i}$ consists of two parts, $a_{i}$ and $b_{i}$.
+$a_{i}$ is the average of L2 norms of all embeddings in the vector space $i$; it represents how important a given head is: the larger the norms, the more attention was given to this attention head.
+$b_{i}$ is the average of cosine distances between all (or a randomly sampled subset, if the user wants to reduce pre-compute time) embeddings in vector space $i$.
 Algorithm 1 Importance scores for heads.
@@
   end for
-Algorithm 1 details the construction of importance scores.
-It is a heuristic based on extensive empirical evaluation; it gives high-quality results across the tested datasets and tasks.
-Intuitively, the score $s_{i}$ of a given head $h_{i}$ consists of two parts, $a_{i}$ and $b_{i}$.
-$a_{i}$ is the average of L2 norms of all embeddings in the vector space $i$; it represents how important a given head is: the larger the norms, the more attention was given to this attention head.
-$b_{i}$ is the average of cosine distances between all (or a randomly sampled subset, if the user wants to reduce pre-compute time) embeddings in vector space $i$.
 Intuitively, $b_{i}$ is a proxy for measuring the “spread” of vector space $i$: the larger $b_{i}$, the larger the average angle between different embeddings in this space is.
@@
 Finally, all chunks from all lists are sorted using their weights and the top $k$ chunks form the final list.
-2.3.1 Integration with Data Stores
+2.3.1 Integration with Data Stores & Other Types of Models
 MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches.
@@
 These two parts of the broader RAG processing pipeline are orthogonal to MRAG.
+Similarly, MRAG does not depend specifically on the embedding form, as long as it is based on a model that harnesses multi-head attention any such approach that results in a valid embedding can be used.
+As such, it could also be used with models such as RetroMAE (Xiao et al., 2022) and the classic BGE-embeddings (Xiao et al., 2022; Chen et al., 2024a).
 3 Multi-Aspect Datasets, Queries, and Metrics
-To assess how well MRAG performs on multi-aspect queries, and to compare it to modern RAG schemes, we need (1) datasets of documents that capture multi-aspectuality, (2) queries to the LLM that touch upon multi-aspectuality and require retrieving documents from the multi-aspect dataset, and (3) metrics that assess how well a RAG scheme retrieves such multi-aspect data.
+To assess how well MRAG performs on multi-aspect queries, and to compare it to modern RAG schemes, we need (1) datasets of documents that capture multi-aspectuality, (2) queries to the LLM that touch upon multi-aspectuality and require retrieving different documents from the multi-aspect dataset, and (3) metrics that asses how well a given RAG scheme retrieves such multi-aspect data.
 We now describe these three elements.
-In Section 4, we also briefly discuss real-world data and queries used.
-Multi-Aspect Dataset Generation We first select conceptually different categories of documents.
-We primarily focus on publicly available Wikipedia articles and select 25 categories (e.g., countries, board games, historical swords, shipwrecks, etc.).
-For each category, we sample 50 documents.
+In Section 4, we also discuss real-world data and queries used to evaluate MRAG.
+3.1 Multi-Aspect Datasets
+We first select conceptually different categories of documents for a synthetic dataset.
+Here, we harness publicly available Wikipedia articles.
+In the dataset construction pipeline, the user selects a given number of categories (e.g., countries, board games, historical swords, shipwrecks, etc.) and then, for each category, they sample a specified number of documents.
 The first part of the document (overview) is used as a text chunk to be embedded.
 We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes.
-Multi-Aspect Query Generation We also require queries that touch upon a given number of $n$ aspects.
+We also use multi-aspect real-world inspired datasets consisting of NDAs and reports describing industry accidents in chemical processing plants.
+We ensure the usefulness of these datasets by working directly with tech leaders from 3 corporations that rely on RAG in their in-house LLM-driven report generation and analytics frameworks.
+Example categories of the legal documents are legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.).
+Examples of accident causes are natural disasters, human mistakes, or lack of proper training.
+We fully release these datasets to propel RAG research.
+Details on all three datasets can be found in the Appendix B.2.
+In our evaluation, we use a total of 13,750 documents.
+3.2 Multi-Aspect Query Generation
+We also require queries that touch upon a given number of $n$ aspects.
 For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories.
@@
 An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 3.
-Metrics We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality.
+3.3 Metrics
+We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality.
 For a query $Q$, a used retrieval strategy $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$.
@@
 Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category.
+While less desired, it still increases chances for a more accurate LLM answer following the retrieval.
+For example, when asking the LLM to determine the cause of an industry accident, fetching the documents in the same category as the accident being queried about, improves the chances for the LLM to give a more relevant answer.
 To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$.
@@
 An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 3.
+Figure 3: An example query used to evaluate different RAG strategies.
+We mention the documents to be fetched in the text and then assess the success ratio of different RAG strategies in finding these documents and their categories.
+We mark exact document matches Refer to caption , category matches Refer to caption , documents that match a category multiple times Refer to caption , and text segments with no matching document Refer to caption .
+Finally, we show the weighted success ratio for each strategy, taking a 2:1 weighting (prioritizing the exact article matches).
 4 Evaluation
 We now illustrate the advantages of MRAG over the state of the art.
-Comparison Baselines We compare MRAG to two main baselines: Standard RAG and Split RAG.
+Comparison Baselines We consider three main baselines: Standard RAG, Split RAG, and Fusion RAG (Rackauckas, 2024).
 The first represents a modern RAG pipeline in which each document uses the activations of the last decoder layer as its embedding.
@@
 The purpose of Split RAG is to show that MRAG’s benefits come from using the multi-head output as embedding and not merely using multiple embedding spaces.
-Additionally, we consider Fusion RAG [30], an optional mechanism that we harness to further enhance the benefits of MRAG at the cost of additional tokens (detailed in Section 4.2).
+Additionally, we consider Fusion RAG (Rackauckas, 2024), an optional mechanism that we harness to further enhance the benefits of MRAG at the cost of additional tokens (detailed in Section 4.3).
 We use queries and metrics introduced in Section 3.
@@
 Among these $k$ documents, we search for matches with the ground truth.
-Figure 3: An example query used to evaluate different RAG strategies.
-We mention the documents to be fetched in the text and then assess the success ratio of different RAG strategies in finding these documents and their categories.
-We mark exact document matches Refer to caption , category matches Refer to caption , documents that match a category multiple times Refer to caption , and text segments with no matching document Refer to caption .
-Finally, we show the weighted success ratio for each strategy, taking a 2:1 weighting (prioritizing the exact article matches).
 Samples & Summaries Each data point in our plots corresponds to 25 queries.
@@
 Our primary focus is on the average retrieval performance among those 25 queries.
-4.1 Analysis of Results
+Figure 4: Retrieval success ratio over 25 queries between MRAG and Standard RAG; each query uses 10 different aspects.
+The top part presents exact document matches, the bottom part presents category only matches (we explain the metrics in Section 3).
+A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (the number of documents fetched for each query is 30).
+Figure 5: Relative retrieval improvement of MRAG over Standard RAG across queries with different numbers of aspects and different embedding models (SFR in the left side, e5 in the right side).
+4.1 Superior Performance for Multi-Aspect Queries
 We start from the query example in Figure 3 and show first the absolute retrieval performance of MRAG over Standard RAG in Figure 4.
@@
 MRAG consistently outperforms the Standard RAG by 10-20% on average, not only across the number of documents fetched, but also across the number of aspects present in the replies, for both models.
-Figure 4: Retrieval success ratio over 25 queries between MRAG and Standard RAG, where each query includes 10 different aspects.
-The upper part presents exact document matches while the lower part presents category only matches (we explain the metrics used in Section 3).
-A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (the number of documents fetched for each query is 30).
-Figure 5: Relative retrieval improvement of MRAG over Standard RAG across queries with different numbers of aspects and different embedding models (SFR in the left side, e5 in the right side).
+4.2 Comparable Performance for Single-Aspect Queries
+We additionally show in Table 1 that MRAG performs on-par with Standard RAG on queries from our multi-aspect dataset where only a single aspect is expected.
+Hence, our approach does not suffer from significant decrease in performance for single-aspect tasks.
+Table 1: Retrieval success ratio (the exact document match) for 25 queries with a single aspect.
 […]
@@
 SFR e5 SFR SFR
-MRAG Standard RAG | MRAG Standard RAG | MRAG Standard RAG | MRAG Standard RAG
+MRAG Standard RAG MRAG Standard RAG MRAG Standard RAG MRAG Standard RAG
 1 24/25 25/25 24/25 25/25 24/25 24/25 25/25 25/25
@@
 3 25/25 25/25 25/25 25/25 25/25 25/25 25/25 25/25
-Table 1: Retrieval success ratio (the exact document match) for 25 queries with a single aspect.
-We additionally show in Table 1 that MRAG performs on-par with Standard RAG on queries from our multi-aspect dataset where only a single aspect is expected.
-Hence, our approach does not suffer from significant decrease in performance for single-aspect tasks.
 4.3 Further Improvements with Additional Tokens
@@
 MRAG performs better than Split RAG, illustrating that its high accuracy is due to the actual multi-head part, and not merely just partitioning the vector and using multiple embedding spaces.
-Figure 6: Relative retrieval improvements of MRAG over Standard RAG for the SFR embedding model compared with Split RAG (the blue plots), and the relative retrieval improvements of Fusion MRAG over both Fusion RAG and MRAG (the red plots).
+Figure 6: Relative retrieval improvements of MRAG over Standard RAG for the SFR embedding model compared with Split RAG (the blue plots), and the relative retrieval improvements of Fusion MRAG over Standard RAG compared with Fusion RAG (the red plots).
 4.5 Real-World Workloads
@@
 4.6 Additional Analyses
-Figure 8: Evaluation of different voting strategies for MRAG and Split RAG
-We also analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one) and the impact of using different voting strategies for MRAG as well as for Split RAG.
-We consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention.
+Figure 8: Evaluation of different voting strategies for MRAG and Split RAG.
+We also analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one).
+Here, we consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention.
 We discover that the last multi-head attention performs the best when compared with the Standard RAG.
-We also illustrate selected representative data from a long investigation two additional voting strategies for MRAG.
+We also illustrate selected representative data from a long investigation into two additional voting strategies for MRAG.
 We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$.
@@
 Figure 8 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
+The complexity of the importance score calculation (Algorithm 1) is $\mathcal{O}(n^{2})$ where $n$ is the number of the embedded documents; it is dominated by calculating the pair-wise cosine similarity and the calculation of the norm.
+Please note that this step needs to be done only once for each dataset and is not a bottleneck.
+Finally, in addition to the performance evaluation, we also investigated the attention heads of the SFR-Embedding-Mistral model as well as LLaMA-2 7B model (model not fine-tuned for text-embedding tasks).
+This analysis is presented in Appendix C.
 5 Related Work
 Our work touches on many areas which we now briefly discuss.
-Many RAG schemes appeared recently [11], using the output of the last decoder layer for embedding generation.
+Many RAG schemes appeared recently (Gao et al., 2024), using the output of the last decoder layer for embedding generation.
 In contrast, MRAG leverages different embedding spaces of attention heads to focus on different aspects of documents and queries.
 As such, it can be combined with other schemes to further improve RAG pipelines.
-Retrieval is sometimes enhanced by a cross-encoder reranking phase [31, 27, 28, 20, 10, 23].
+Retrieval is sometimes enhanced by a cross-encoder reranking phase (Rosa et al., 2022; Nogueira & Cho, 2020; Nogueira et al., 2020; Li et al., 2021; Gao et al., 2021; MacAvaney et al., 2019).
 In such solutions, typically after retrieving a set of relevant chunks, they are re-ranked using specialized models.
@@
 Structure-enhanced RAG schemes employ different strategies for structuring text to improve retrieval quality.
-A common idea is to construct a Knowledge Graph from text, which enables retrieval amongst entities and relationships [18, 7, 17, 4, 38].
-RAPTOR [32] generates multi-level summaries for clusters of related chunks, building a tree of summaries with increasing levels of abstraction to better capture the meaning of the text.
-Graph RAG [8] creates a Knowledge Graph, and summarizes communities in the graph, which provide data at the different levels of abstraction.
+A common idea is to construct a Knowledge Graph from text, which enables retrieval amongst entities and relationships (Jiang et al., 2024; Delile et al., 2024; Hussien et al., 2024; Bui et al., 2024; Xu et al., 2024a).
+RAPTOR (Sarthi et al., 2024) generates multi-level summaries for clusters of related chunks, building a tree of summaries with increasing levels of abstraction to better capture the meaning of the text.
+Graph RAG (Edge et al., 2024) creates a Knowledge Graph, and summarizes communities in the graph, which provide data at the different levels of abstraction.
 All these systems try to improve RAG quality by utilizing additional structures that describe entity relationships or the inner organization of text.
@@
 By addressing the challenges of multi-aspectuality in queries, MRAG paves the way for more reliable and accurate LLM applications across diverse industries.
-Acknowledgments and Disclosure of Funding
+Acknowledgements
 We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines, and for their excellent technical support.
@@
 B.2 Dataset Details
+Table 2: Overview of the structure and the number of documents in the respective datasets.
+dataset #categories #topics #documents total #documents
+Wikipedia 25 50 documents per category 1250
+Legal Documents 25 25 per category 10 per topic 6250
+Accident Reports 25 25 per category 10 per topic 6250
+B.3 Prompt Template for the Synthetic Dataset Generation
 Table 3: Prompt template for query generation.
@@
 Again, make sure that you reference all the following topics in your story: <list of titles>
+Appendix C Attention Head Analysis
+We investigated the attention heads of two models in detail: LLaMA-2 7B and SFR-Embedding-Mistral.
+We selected these two models for a detailed investigation because the former represents models that are not fine-tuned for text embeddings, while the latter is specifically the text embedding model that we used for our experiments.
+For each model, we looked specifically at the attention scores within each attention head, i.e., how much attention each head pays to each input token during the inference.
+Knowing the semantics of the input tokens enables then deriving certain conclusions about multi-aspectuality and attention heads.
+We plot selected results in Figure 9.
+Each heatmap shows the dot-product between key- and value-projections inside a given specified attention head, where line $i$ of a heatmap for attention head h indicates the dot-products between the query-projection of token $i$ and the key-projections of all previous tokens $j<i$ (both models use causal attention).
+(a) Head 4 LLaMA-2
+(b) Head 8 SFR
+(c) Head 22 LLaMA-2
+(d) Head 21 SFR
+Figure 9: Heatmap plots for selected attention heads of the LLaMA-2 7B and SFR-Embedding-Mistral models.
+For both models, we found out that the attention patterns vary significantly between the different attention heads.
+Still, we encountered two distinct patterns.
+First, the diagonal lines in Figures 9(a) and 9(b) indicate that, when processing a certain input token x, elevated attention is paid to some tokens that came a constant numbers of steps before x.
+We postulate that this pattern is likely beneficial to understanding the overall rhythm of a natural language, allowing the model to better identify which words are semantically connected, and which parts of the input text refer to each other.
+Second, horizontal and vertical lines in Figures 9(c) and 9(d) show that these heads learned to pay attention to specific tokens, regardless of how far apart they are within the input sequence.
+An intuitive justification for such patterns is the focus on certain semantic aspects of the input sequence.
+(a) Attention Head 0 LLaMA-2 7B
+(b) Attention Head 8 SFR
+(c) Attention Head 14 LLaMA-2 7B
+(d) Attention Head 21 SFR
+Figure 10: Attention scores for selected attention heads of the LLaMA-2 7B and SFR-Embedding-Mistral models.
+We also detail attention scores (after applying softmax) of selected heads in Figures 10 and 11, when the model is processing the last token of its input.
+We see that some tokens gather a lot of attention from most heads, yet there is always a plethora of passages which are attended differently by any two attention heads.
+An interesting pattern we encountered was that for the SFR-Embedding-Mistral model (see Figure 11), all heads’ attention spiked significantly on the first line-break in the input sequence - either positively or negatively.
+We conjecture that this is a consequence of how the embedding model was fine-tuned and its intended usage pattern: embedding queries are usually prepended with a retrieval instruction, which is terminated by a line-break.
+The model likely learnt to summarise the necessary information about this instruction inside the terminating line-break.
+Figure 11: Attention scores for all attention heads of the SFR-Embedding-Mistral model.

### v5

@@
 This paper introduces Multi-Head RAG (MRAG), a novel scheme designed to address this gap with a simple yet powerful idea: leveraging activations of Transformer’s multi-head attention layer, instead of the decoder layer, as keys for fetching multi-aspect documents.
-The driving motivation is that different attention heads can learn to capture different data aspects.
+The driving observation is that different attention heads learn to capture different data aspects.
 Harnessing the corresponding activations results in embeddings that represent various facets of data items and queries, improving the retrieval accuracy for complex queries.
-We provide an evaluation methodology and metrics, multi-aspect datasets that we release online, and real-world use cases to demonstrate MRAG’s effectiveness, showing improvements of up to 20% in relevance over standard RAG baselines.
-MRAG can be seamlessly integrated with existing RAG frameworks and benchmarking tools like RAGAS as well as different classes of data stores.
+We provide an evaluation methodology and metrics, multi-aspect datasets, and real-world use cases to demonstrate MRAG’s effectiveness.
+We show MRAG’s design advantages over 18 RAG baselines, empirical improvements of up to 20% in retrieval success ratios, and benefits for downstream LLM generation.
+MRAG can be seamlessly integrated with existing RAG frameworks and benchmarks.
 1 Introduction
-Large Language Models (LLMs) transformed many machine learning tasks using in-context learning abilities.
-They achieved such accuracy by leveraging an increasing number of parameters, which in recent models have grown to hundreds of billions, making LLM training expensive in terms of both time and resources.
-It also comes with the danger of leaking confidential data into model weights (Yan et al., 2024; Shi et al., 2024; Patil et al., 2024).
-Additionally, continuous training through fine-tuning is necessary to keep LLMs up-to-date.
-Even using the newest data, LLMs display an ongoing problem of hallucinations (Zhang et al., 2023; Xu et al., 2024c; Huang et al., 2023) by providing factually incorrect information.
-Retrieval Augmented Generation (RAG) was proposed (Lewis et al., 2020; Guu et al., 2020) in order to address these issues as well as others and make LLMs more trustworthy.
-The key idea behind RAG is to enhance the generative model’s capabilities by integrating a retrieval system that fetches relevant passages from a large corpus of data.
-In this setting, when a query is received, the retrieval system first identifies and retrieves pertinent information, which is fed into the generative model’s context for a more accurate and relevant response.
-Instead of the model storing information within its weights, RAG effectively leverages external knowledge, reducing hallucinations (by grounding the LLM reply in reliable sources), and ensuring that responses contain up-to-date knowledge (e.g., by accessing the Internet), all without requiring expensive training.
-More specifically, there are two main stages in a RAG pipeline: data preparation and query execution.
-During data preparation, one constructs a vector database (DB) populated with embeddings and their corresponding data items such as documents.
-During query execution, one constructs an embedding of that query and retrieves data items in the store with similar embeddings.
-Intense recent research efforts have been put into RAG (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2024; Huang & Huang, 2024; Yu et al., 2024; Mialon et al., 2023; Li et al., 2022).
-On one hand, different RAG designs have been proposed, for example RAPTOR (Sarthi et al., 2024), Self-RAG (Asai et al., 2023), Chain-of-Note (Yu et al., 2023), and many others (Abdallah & Jatowt, 2024; Delile et al., 2024; Edge et al., 2024; Manathunga & Illangasekara, 2023; Zeng et al., 2024; Wewer et al., 2021; Xu et al., 2024b).
-In general, these schemes focus on making the retrieved data more accurate and relevant to the query.
-There have also been efforts into benchmarking and datasets for RAG evaluation (Chen et al., 2024b; Xiong et al., 2024; Lyu et al., 2024; Es et al., 2023).
-Despite all these advances, we observe that no existing RAG scheme or evaluation methodology explicitly targets an important class of problems that come with a high degree of multi-aspectuality.
-These are problems that require combining several (potentially many) significantly different aspects in a single query.
-As a simple illustrative example of such a query, consider the question “What car did Alexander the Great drive?”, and assume that the queried model has not been trained on history.
-When using RAG, to answer this question accurately, one would retrieve two documents, one describing Alexander the Great and one outlining the history of car manufacturing.
-However, the embeddings of these two documents could be far away from each other in the embedding space.
-At the same time, such queries are common in different industry settings, as indicated by extensive discussions with our industry collaborators.
-Imagine a chemical processing plant experiencing an equipment accident.
-One could use an LLM to find the accident cause, which might require the retrieval of multiple, potentially confidential documents to provide the necessary context.
-These documents could be related to different aspects, for example psychological profiles of workers (“Was the accident due to mismanaging a worker?”), equipment purchase records (“Was some equipment part too old?”), maintenance (“Was some equipment part rusty?”), weather (“Was there a particularly strong thunderstorm at the accident time that could have caused dangerous power spikes in the grid?”), or even microclimate (“Was it too humid for an extended period of time in the production hall?”).
-As we illustrate in Section 4, such problems pose challenges for existing RAG schemes and have been unaddressed by modern RAG benchmarking pipelines.
+Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020) emerged as a promising remedy for several key limitations of Large Language Models (LLMs).
+By decoupling knowledge from model weights, RAG reduces the risk of leaking confidential data (Yan et al., 2025; Shi et al., 2024; Patil et al., 2024), a critical concern when training on sensitive corpora.
+It also mitigates hallucinations (Zhang et al., 2025; Xu et al., 2025; Huang et al., 2025b) by grounding LLM outputs in retrieved, verifiable information.
+The core mechanism involves augmenting a generative LLM with a retrieval module that fetches relevant passages from an external corpus in response to a query.
+Rather than relying solely on static, parametric knowledge, RAG dynamically incorporates retrieved content into the model’s context, enabling more accurate and up-to-date responses.
+While early RAG systems required training task-specific retrievers and readers (Humeau et al., 2020; Singh et al., 2021), the current trend favors lightweight, in-context learning (ICL) approaches (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2025; Huang & Huang, 2024), which avoid the cost and complexity of retraining and allow for rapid knowledge updates without modifying the underlying LLM parameters.
+A RAG pipeline consists of two main stages: data preparation and query execution.
+In the preparation stage, a vector database (DB) is constructed by embedding a collection of documents and storing these embeddings alongside their associated content.
+At inference time, the query is similarly embedded, and nearest-neighbor search retrieves the most relevant data items, which are then passed to the LLM for final answer generation.
+Ongoing developments in RAG have led to variety of different RAG designs (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2025; Huang & Huang, 2024; Yu et al., 2024a; Mialon et al., 2023; Li et al., 2022; Chen et al., 2024c; Xiong et al., 2024; Lyu et al., 2025; Es et al., 2024; Sarthi et al., 2024; Asai et al., 2024; Abdallah & Jatowt, 2024; Delile et al., 2024; Edge et al., 2025; Manathunga & Illangasekara, 2023; Zeng et al., 2024; Wewer et al., 2021; Xu et al., 2024; Huang et al., 2025a; Yang et al., 2025; Su et al., 2025; Zhao et al., 2025a; b; Chen et al., 2024d; Chan et al., 2024; Jiang et al., 2024; Yu et al., 2024b).
+Yet, no existing RAG method or benchmark explicitly targets multi-aspectual problems, that is, queries requiring the integration of multiple, semantically distinct aspects.
+For example, answering “What car did Alexander the Great drive?” (assuming no historical pretraining) requires retrieving unrelated documents on Alexander the Great and on car manufacturing, whose embeddings may lie far apart in the vector space.
+Such multi-aspect queries are common in industrial settings, as confirmed by extensive discussions with our industry collaborators and further supported by our analysis of over 35 industry reports (Office of Railroad Safety, 2014; Columbia Accident Investigation Board, 2003; Zografos et al., 2013; Reason et al., 2006; Kalia, 2013; Packham, 2017; Messiou et al., 2023; Multer et al., 2013; Coury et al., 2010; Harle, 1997; O’Hare, 2000; Gordon et al., 2005; Bridger, 2021; OpenText, 2022) (details are in Appendix A; we considered accident prevention, healthcare, airport management, and others).
+For example, in a chemical plant accident, determining the cause might require accessing diverse and confidential documents related to worker psychology (“Was it mismanagement?”), equipment records (“Was a part outdated or rusty?”), weather conditions (“Were there power spikes due to a storm?”), or even microclimate (“Was prolonged humidity a factor?”).
+As shown in Section 5, such cases have been unaddressed by modern RAG schemes and benchmarks (Chen et al., 2024c; Xiong et al., 2024; Lyu et al., 2025; Es et al., 2024).
 Figure 1: An overview of the decoder architecture, and a comparison of how standard RAG and Multi-Head RAG embeddings are generated.
-In this work, we propose Multi-Head RAG (MRAG): a scheme that addresses the above problem.
-Common practice in modern RAG designs is the use of embeddings based on last-layer decoder block activations.
-Our key idea is to use instead the activations of the multi-head attention part of the decoder block as embeddings.
-The Transformer architecture can be seen as a pipeline with many (e.g., 96 for GPT-3 (Wang et al., 2024b)) blocks, where a single block consists of an attention module and a feed-forward module.
-Each individual attention module is multi-headed: it consists of multiple parts called heads that learn different sets of weight matrices; see Figure 1 for an overview.
-It is conjectured that these different heads could capture different aspects of the processed data.
-We use this as a driving design feature that facilitates capturing the potential multi-aspectuality of the data without increasing space requirements compared to standard RAG, and without any fine-tuning or other modifications to the harnessed model (contribution 1).
-Such multi-aspect embeddings are then directly used for both data items and query representation.
-Considering multi-aspectuality explicitly comes with challenges.
+In this work, we propose Multi-Head RAG (MRAG): a scheme that addresses the above problem (contribution 1).
+Common practice in many modern RAG designs is to use embeddings derived from last-layer decoder block activations of a decoder-based embedding LLM (a language model specifically fine-tuned to provide high-quality embeddings).
+Examples of such embedding models are SFR-Embedding-Mistral and E5-Mistral-7B.
+Our key idea is to extend this design by incorporating activations from the multi-head attention (MHA) modules of decoder blocks as embedding sources.
+This enables the representation of multiple distinct aspects of the input text.
+Specifically, a Transformer consists of a stack of blocks (e.g., 96 in GPT-3 (Wang et al., 2025b)), each containing an MHA module with multiple heads that are trained with separate parameter sets.
+Through a survey of literature into Transformer design and interpretability, we find empirical and theoretical support for the conjecture that different heads specialize in different aspects of the input (details are in Section 2.4).
+This enables efficient multi-aspect embeddings without increasing space or compute costs compared to standard RAG, and without requiring any additional fine-tuning or architectural modifications to the base model.
+Considering multi-aspectuality comes with challenges.
 For example, it is unclear how to assess whether a RAG solution does indeed harness multiple aspects when fetching documents.
-For this, we establish an evaluation methodology as well as a full data construction and query processing pipeline that implements the multi-aspect embedding idea (contribution 2).
-Our datasets facilitate broad evaluation by considering both fully-automatically generated, synthetic data as well as data based on specific industry use cases that show the benefits of MRAG (contribution 3).
+For this, we develop a multi-aspect RAG pipeline that includes data preparation and query processing with both multi-aspect retrieval and ranking schemes (contribution 2).
+We also establish an evaluation methodology and provide multi-aspect datasets, complementing existing RAG benchmarks (Chen et al., 2024c; Xiong et al., 2024; Lyu et al., 2025; Es et al., 2024) (contribution 3).
 We ensure the relevance of our RAG datasets in real use cases by working directly with tech leaders (e.g., a generative AI division head) from 3 corporations, all of which actively use RAG in their own LLM infrastructures.
-Our evaluation illustrates the benefits in the relevance of retrieved documents, for example 20% over a modern RAG baseline for fetching multi-aspect Wikipedia articles, and comparable performance for single-aspect queries (contribution 4).
-We also show how MRAG and its benchmarking principles can be seamlessly integrated with both existing RAG solutions and benchmarking frameworks such as RAGAS (contribution 5).
+We illustrate the advantages of MRAG over 18 traditional and modern RAG designs in various design criteria and in both time and space complexities (contribution 4).
+In evaluation, MRAG enhances the relevance of retrieved documents by up to 20% over modern RAG baselines, offers comparable performance without degradation for single-aspect queries, and benefits the downstream LLM generation (contribution 5).
+Thanks to its simplicity, MRAG can be seamlessly integrated into any stores while its benchmarking methodology can straightforwardly extend benchmarks such as RAGAs (contribution 6).
 MRAG’s code is publicly available [^1] https://github.com/spcl/MRAG.
-2 The MRAG Formulation & Pipeline
-We now present in detail the mathematical underpinning of MRAG and its corresponding pipeline.
-Decoder Formulation We first introduce formally the decoder architecture.
-We omit, for clarity, unnecessary details such as layer normalizations.
-The input is a text chunk that consists of $n$ tokens.
-The output of attention head $h$ for the $i$th token $x_{i}$ is defined as follows (Vaswani et al., 2017):
-$\text{head}^{h}(\mathbf{x}_{i})=\sum_{j}w_{ij}\mathbf{v}_{j}^{h},\quad\text{where}$ (1)
-$\vspace{-0.5em}w_{ij}=\text{softmax}\left(\left(\mathbf{q}_{i}^{h}\right)^{T}\mathbf{k}_{j}^{h}\right),\quad\mathbf{q}_{i}^{h}=\mathbf{W}_{q}^{h}\mathbf{x}_{i},\quad\mathbf{k}_{j}^{h}=\mathbf{W}_{k}^{h}\mathbf{x}_{j},\quad\mathbf{v}_{j}^{h}=\mathbf{W}_{v}^{h}\mathbf{x}_{j}$ (2)
-where $\mathbf{W}_{q}^{h},\mathbf{W}_{k}^{h},\mathbf{W}_{v}^{h}$ are, respectively, learnable query, key, and value projections associated with head $h$, and $\mathbf{x}_{j}$ is the vector embedding of the $j$th token $x_{j}$.
-These outputs get combined to form the output of the $i$th multi-head attention block as follows:
-[…]
-where matrix $\mathbf{W}_{o}$ is the linear layer that combines the outcomes of all the attention heads. This step is then followed by the Transformer feed-forward layer.
-[…]
-Multi-Head RAG Formulation The key idea behind MRAG is simple: instead of the single activation vector generated by the last feed-forward decoder layer for the last token, we harness the $H$ separate activation vectors generated by the last attention layer for the last token, before processing it via $\mathbf{W_{o}}$.
-This can be formulated as a set of embeddings $\mathcal{S}$:
-$\mathcal{S}=\{\mathbf{e}_{k}\forall_{k}\}\quad\text{where}\quad\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})$ (4)
-which is simply the set of all outputs from the attention heads on the last token $\mathbf{x}_{n}$ of the input.
-As processing with multiple heads does not change the size of the output vector, $\mathcal{S}$ has the same space requirements as standard RAG.
-However, because we capture the separate embeddings before their mixing with $\mathbf{W}_{o}$, we conjecture that it gives more information about what the different parts of the input attend to, facilitating capturing multi-aspectuality.
-Naming We use the terms “single-aspect embedding” and “multi-aspect embedding” to refer to, respectively, a small embedding extracted from a single attention head and a collection of all single-aspect embeddings extracted from an attention layer.
-2.1 Overview of the Multi-Head RAG Pipeline
-We now describe how the above embedding model fits the RAG pipeline.
-Figure 2 shows a summary of the design.
-The MRAG pipeline consists of two main parts, dedicated to data preparation [Uncaptioned image] and query execution [Uncaptioned image].
-Both parts heavily use the data store [Uncaptioned image] (vector DB).
-2.1.1 Data Preparation
-When preparing data [Uncaptioned image], we populate a data store [Uncaptioned image] with multi-aspect MRAG text embeddings [Uncaptioned image] and their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image] (MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data).
-We create the multi-aspect embedding [Uncaptioned image] of each text chunk [Uncaptioned image] using a selected decoder-based embedding model [Uncaptioned image] (this part is detailed in Section 2.2).
-The user of the pipeline can plug in their model [Uncaptioned image] of choice as well as use their input data.
-We also offer a dedicated synthetic data generator [Uncaptioned image] that can be used to construct multi-aspect input documents [Uncaptioned image] (we detail this part in Section 3) for evaluation purposes.
-MRAG stores data differently than standard RAG, where a single embedding [Uncaptioned image] points to a single text chunk [Uncaptioned image].
-For MRAG, each multi-aspect embedding consists of $h$ single-aspect embeddings [Uncaptioned image], each pointing to the original text chunk [Uncaptioned image].
-So the data store [Uncaptioned image] contains $h$ embedding spaces, each capturing a different aspect of the text.
-This crucial feature allows MRAG to compare query [Uncaptioned image] and text chunks [Uncaptioned image] in multiple embedding spaces that capture multiple aspects of the data.
-2.1.2 Query Execution
-During query execution [Uncaptioned image], we first generate a multi-aspect embedding [Uncaptioned image] of the input query [Uncaptioned image], using the selected embedding model [Uncaptioned image] (details in Section 2.2).
-Then, we find the nearest multi-aspect embeddings [Uncaptioned image] and their corresponding text chunks [Uncaptioned image] in the data store [Uncaptioned image] using a special multi-aspect retrieval strategy [Uncaptioned image] (detailed in Section 2.3).
-We ensure that there is no overhead in latency due to multiple aspects because computing these different smaller embeddings is done fully in parallel.
-Finally, the retrieved data can optionally be assessed [Uncaptioned image] with novel metrics regarding how well it corresponds to the multi-aspect requirements (detailed in Section 3).
-As with the data preparation [Uncaptioned image] stage, the query execution [Uncaptioned image] stage is flexible, and the user can plug in their models [Uncaptioned image] / [Uncaptioned image] of choice and use their own queries [Uncaptioned image].
-We also offer a dedicated synthetic query generator [Uncaptioned image] that can be used to construct multi-aspect input queries [Uncaptioned image] (detailed in Section 3) for evaluation purposes.
+2 MRAG: Design & Implementation
+A typical RAG scheme (see Figure 2) consists of two main parts: data preprocessing [Uncaptioned image] and query execution [Uncaptioned image]; both parts heavily use an embedding model [Uncaptioned image] and the data store [Uncaptioned image].
+During preprocessing, each document in the database is encoded into one or more embeddings using the embedding model; these embeddings are stored with that document (usually as a key-value pair in a vector DB).
+During query execution, the embedding of the user-provided query is constructed using the same embedding model; then, the retriever fetches candidate documents based on embedding similarity, the reranker refines their ordering using more precise scoring, and the reader uses a downstream generative model to synthesize the final answer from the top-ranked documents.
 Figure 2: Overview of the MRAG pipeline, consisting of two parts: data preparation Refer to caption and query execution Refer to caption.
@@
 2.1 Constructing Multi-Aspect Embeddings [Uncaptioned image]
-MRAG can leverage any embedding model with multi-head attention support to construct the multi-aspect embeddings for a given input text.
-In this work, we consider two embedding models from the MTEB leaderboard (Huggingface, 2024) as potential candidates.
-Specifically, the SFR-Embedding-Model (Meng et al., 2024) and the e5-mistral-7b-instruct (Wang et al., 2024a), both based on the Mistral 7B architecture with 32 decoder blocks and 32 attention heads per multi-head attention.
-While our approach allows for extracting and using the multi-aspect embeddings from any decoder block, and from different layers within a block, we found that multi-aspect embeddings extracted from the last multi-head attention worked best in our experimental setting.
-We provide further discussion on the carried out experiments in Section 4.
-2.3 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
-A retrieval strategy determines how we select the closest text chunks from the DB given a multi-aspect embedding of the user query.
-In general, the MRAG retrieval strategy consists of three steps.
-First, during data preparation, we assign importance scores to all $h$ embedding spaces.
-Intuitively, these scores capture the fact that different spaces (and the corresponding heads) may be more or less relevant for the used data.
-Then, during query execution, MRAG starts by applying the traditional RAG retrieval separately for each embedding space.
-This returns a list of $c$ closest text chunks for each embedding space (a total of $h$ lists).
-Here, we use a special voting strategy to pick overall top $k$ out of all $hc$ chunks, using the pre-computed importance scores.
-Algorithm 1 details the construction of importance scores.
-It is a heuristic based on extensive empirical evaluation; it gives high-quality results across the tested datasets and tasks.
-Intuitively, the score $s_{i}$ of a given head $h_{i}$ consists of two parts, $a_{i}$ and $b_{i}$.
-$a_{i}$ is the average of L2 norms of all embeddings in the vector space $i$; it represents how important a given head is: the larger the norms, the more attention was given to this attention head.
-$b_{i}$ is the average of cosine distances between all (or a randomly sampled subset, if the user wants to reduce pre-compute time) embeddings in vector space $i$.
-Algorithm 1 Importance scores for heads.
-  for each head $h_{i}$ do
-     $a_{i}\leftarrow 0$; $b_{i}\leftarrow 0$
-     $count\_a_{i}\leftarrow 0$; $count\_b_{i}\leftarrow 0$
-     for each embedding $e_{ij}$ in $h_{i}$ do
-        $a_{i}\leftarrow a_{i}+||e_{ij}||$
-        $count\_a_{i}\leftarrow count\_a_{i}+1$
-        for each embedding $e_{ih}$ do
-           $b_{i}\leftarrow b_{i}+\text{cosine-distance}(e_{ij},e_{ih})$
-           $count\_b_{i}\leftarrow count\_b_{i}+1$
-        end for
-     end for
-     $a_{i}\leftarrow a_{i}/count\_a_{i}$; $b_{i}\leftarrow b_{i}/count\_b_{i}$
-     $s_{i}\leftarrow a_{i}\cdot b_{i}$
-  end for
-Intuitively, $b_{i}$ is a proxy for measuring the “spread” of vector space $i$: the larger $b_{i}$, the larger the average angle between different embeddings in this space is.
-Deriving $s_{i}$ as a product $a_{i}\cdot b_{i}$ ensures that we reward heads with high average attention and high average spread, but simultaneously penalize heads with lower average attention or with low average spread (both $a_{i}$ and $b_{i}$ are appropriately scaled).
-The used voting strategy combines the constructed lists of text chunks from individual embedding spaces into a single list of top $k$ chunks.
-The strategy is very simple (the corresponding Algorithm 2 is in the Appendix).
-Each text chunk from a list $i$ of the vector space $i$ has a certain position on this list, we denote this position with $p$.
-We obtain a weight for this chunk as $s_{i}\cdot 2^{-p}$; $s_{i}$ is the previously defined importance score of the space $i$.
-Multiplying $s_{i}$ with $2^{-p}$ exponentially lowers the significance of less relevant text chunks.
-Finally, all chunks from all lists are sorted using their weights and the top $k$ chunks form the final list.
-2.3.1 Integration with Data Stores & Other Types of Models
-MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches.
-It can be combined with both the exact and the approximate NN to find the matching (embedding, chunk)-pairs.
-These two parts of the broader RAG processing pipeline are orthogonal to MRAG.
-Similarly, MRAG does not depend specifically on the embedding form, as long as it is based on a model that harnesses multi-head attention any such approach that results in a valid embedding can be used.
-As such, it could also be used with models such as RetroMAE (Xiao et al., 2022) and the classic BGE-embeddings (Xiao et al., 2022; Chen et al., 2024a).
-3 Multi-Aspect Datasets, Queries, and Metrics
-To assess how well MRAG performs on multi-aspect queries, and to compare it to modern RAG schemes, we need (1) datasets of documents that capture multi-aspectuality, (2) queries to the LLM that touch upon multi-aspectuality and require retrieving different documents from the multi-aspect dataset, and (3) metrics that asses how well a given RAG scheme retrieves such multi-aspect data.
-We now describe these three elements.
-In Section 4, we also discuss real-world data and queries used to evaluate MRAG.
-3.1 Multi-Aspect Datasets
-We first select conceptually different categories of documents for a synthetic dataset.
-Here, we harness publicly available Wikipedia articles.
-In the dataset construction pipeline, the user selects a given number of categories (e.g., countries, board games, historical swords, shipwrecks, etc.) and then, for each category, they sample a specified number of documents.
-The first part of the document (overview) is used as a text chunk to be embedded.
-We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes.
-We also use multi-aspect real-world inspired datasets consisting of NDAs and reports describing industry accidents in chemical processing plants.
-We ensure the usefulness of these datasets by working directly with tech leaders from 3 corporations that rely on RAG in their in-house LLM-driven report generation and analytics frameworks.
-Example categories of the legal documents are legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.).
-Examples of accident causes are natural disasters, human mistakes, or lack of proper training.
-We fully release these datasets to propel RAG research.
-Details on all three datasets can be found in the Appendix B.2.
-In our evaluation, we use a total of 13,750 documents.
-3.2 Multi-Aspect Query Generation
-We also require queries that touch upon a given number of $n$ aspects.
+An embedding is constructed for each data item in a pre-existing database (part [Uncaptioned image]) and for the user query (part [Uncaptioned image]).
+In standard modern RAG, given an input chunk of $n$ tokens constituting a document or chunk to be embedded, one obtains a corresponding embedding by applying the embedding model to the data item and extracting the activation of the final feed-forward layer for the last token $\mathbf{x}_{n}$, i.e., $\texttt{feed-forward}(\texttt{multi-head}(\mathbf{x}_{n}))\in\mathbb{R}^{d}$.
+Appendix B provides additional mathematical details.
+In MRAG, instead of relying on this single embedding, the activations of all $h$ attention heads are obtained before they are merged by the final projection layer.
+Specifically, for the last token $\mathbf{x}_{n}$, we extract the set of head-specific vectors $\mathcal{S}=\{\mathbf{e}_{k}\}_{k=1}^{h}$, where each $\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})\in\mathbb{R}^{d/h}$ is referred to as a “single-aspect embedding” and $\mathcal{S}$ is called a “multi-aspect embedding”.
+This provides $h$ semantically diverse embeddings per input, reflecting the different perspectives captured by each attention head.
+Since these vectors are extracted from the same internal activation used by standard RAG, the overall space and time requirements remain unchanged.
+However, the explicit separation of heads allows MRAG to preserve fine-grained aspectual distinctions, facilitating multi-aspect representation.
+2.2 Data Preprocessing [Uncaptioned image]
+We populate a data store [Uncaptioned image] with multi-aspect embeddings [Uncaptioned image] for their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image].
+MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data.
+Unlike in standard RAG, where a single embedding [Uncaptioned image] points to a single text chunk [Uncaptioned image], in MRAG, each out of $h$ single-aspect embeddings [Uncaptioned image] points to the original text chunk [Uncaptioned image] (i.e., the data store [Uncaptioned image] contains $h$ embedding spaces).
+This crucial feature allows MRAG to compare query [Uncaptioned image] and text chunks [Uncaptioned image] in multiple embedding spaces that capture multiple aspects of the data.
+MRAG performs one-time importance scoring for each embedding space $i$, capturing its relevance to the dataset.
+Each score $s_{i}$ combines: (i) the average L2 norm $a_{i}$ of embeddings in space $i$, reflecting attention strength, and (ii) the average cosine distance $b_{i}$ between sampled embeddings, approximating its semantic spread.
+The final score is $s_{i}=a_{i}\cdot b_{i}$, encouraging both head relevance (through high $a_{i}$) and high representational diversity (through high $b_{i}$).
+Full details are provided in Algorithm 1, Appendix C.1.
+2.3 Query Execution
+During query execution, MRAG first generates a multi-aspect embedding $\mathcal{S}$ of the input query (cf.
+Section 2.1).
+These embeddings are computed fully in parallel during the same inference pass, so this multi-vector representation introduces no additional latency.
+Retriever.
+Given $\mathcal{S}$, the retriever conducts parallel retrieval of the top-$c$ nearest chunks within each embedding space, and the aggregation of the resulting candidates across all $h$ spaces.
+Reranker.
+Once $hc$ candidate chunks are retrieved across all heads, the ranker stage consolidates them into a single list of top-$k$ results using a simple but effective voting strategy.
+For each candidate chunk at position $p$ in the ranked list for space $i$, we assign a score of $s_{i}\cdot 2^{-p}$, where $s_{i}$ is the precomputed importance score.
+This exponentially discounts lower-ranked candidates and balances influence across heads.
+The final top-$k$ list is obtained by globally sorting all candidates by these scores.
+The voting procedure is described in Algorithm 2 in Appendix C.2.
+Reader.
+The top-$k$ retrieved results are inserted into the LLM context using a multi-aspect prompt template (prompts are fully specified in Appendix D).
+Each result is placed in a separate section of the prompt.
+Stored metadata can be included alongside each chunk to provide additional context, such as the aspect or the chunk category.
+2.4 Capturing Multi-Aspectuality Without Additional Training
+A key design decision in MRAG is to use the hidden representations immediately after the attention block in the last decoder layer, without additional fine-tuning, to avoid training overhead and make deployment easy.
+This is motivated by growing evidence that attention heads in Transformer models naturally converge during training to focus on distinct aspects of the input.
+For example, Wang et al.
+(2025a) show that heads diverge into clusters specialized for different input patterns, while Olsson et al.
+(2022) and others (McDougall et al., 2024; Gould et al., 2024) discover that heads focus on repeated sequences or named entities.
+Similar trends are observed in BERT (Clark et al., 2019; Kovaleva et al., 2019; Htut et al., 2019).
+We provide more details in a brief literature survey (Appendix E.1).
+Given these findings, we assume that even without additional fine-tuning, the embeddings output by MHA already encode multi-aspectuality suitable for downstream retrieval.
+Our own brief attention pattern analysis (Appendix E.2) confirm shifting token focus across heads.
+2.5 Parallelization and Systems Considerations
+MRAG evenly distributes the embedding dimensionality across attention heads, keeping total storage and compute $O(nd)$ (same as standard RAG; we show a more detailed complexity analysis in the following section).
+It uses off-the-shelf ANN indexes that support parallel subspace search, and it avoids dynamic or variable-length vectors that complicate indexing and caching.
+Modern vector databases already support the parallel multi-vector search required by our approach.
+Milvus can execute distributed, parallel ANN queries across shards with low latency (Wang et al., 2025c; Clavié et al., 2024; Wang et al., 2021).
+Pinecone provides cascading retrieval with fixed-size multi-vector encodings (such as ConstBERT) to maintain predictable costs.
+Furthermore, ESPN tackles multi-vector retrieval at SSD and GPU levels by implementing prefetching pipelines and storage bypass mechanisms, achieving near-memory query latency, even when dealing with large indices (Shrestha et al., 2024).
+Taken together, these technologies illustrate that multi-vector retrieval, especially when using fixed-size vectors as MRAG does, can be implemented at scale with negligible practical overhead.
+3 Compute & Storage Complexity Analysis
+We analyze the runtime and space complexity of MRAG and 18 RAG baselines, the results are in Table 1 (the runtime metrics of work/depth and the notation are explained in the table caption).
+Derivation details and a broader discussion are in Appendix F.
+Overall, MRAG achieves competitive results in all aspects.
+At inference time, it extracts $h$ attention-head embeddings in parallel from a single forward pass, resulting in the same latency as standard RAG.
+Preprocessing is lightweight, requiring only one pass per document and simple statistics for head scoring, avoiding the cost of training or complex structure construction.
+Storage overhead is minimal, as $h$ single-aspect embeddings per data item have the same dimensionality as a standard embedding.
+In contrast, prior schemes like Poly-encoder (Humeau et al., 2020) and ColBERT (Khattab & Zaharia, 2020) incur significant cost due to using many token-level embeddings or time-consuming training rounds.
+Table 1: Time and storage complexity of different RAG schemes; MRAG is the only scheme to match the results of a plain vanilla RAG.
+Work is the total number of all operations, Depth is runtime complexity assuming unlimited parallel processing units; both are established measures of analyzing parallel algorithms.
+Retrieval describes steps during inference, including any postprocessing before returning documents or summaries.
+Preprocessing includes all steps before inference.
+$n$: number of data items in a database (i.e., document chunks), $d$: embedding dimensionality, $k$: number of retrieved top chunks per single user query, $l_{q}$: average token length of the query, $l_{d}$: average token length of the document, $W_{m}$$D_{m}$: work/depth to run a transformer-based model, $W_{e}$$D_{e}$: work/depth to embed a graph, $W_{i}$$D_{i}$: work/depth to index a graph, $s$: polynomial function that models the complexity of various additional scheme-specific design decisions, heuristics, etc., which cannot be straightforwardly modeled with closed-form expressions (e.g., count of graph communities, toy-graph size, self-note length, BM25 matching cost, keyword matching cost); it scales at most linearly with $n$ and is typically considerably larger than $O(1)$.
+Scheme Retrieval Preprocessing Storage
+Work Depth Work Depth
+Vanilla RAG $O(W_{m}+nd)$ $O(D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+Poly-encoders (Humeau et al., 2020) $O(W_{m}+nd+s(n)d)$ $O(D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+Lewis et al. (2020) $O(2W_{m}+nd)$ $O(2D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+ColBERT (Khattab & Zaharia, 2020) $O(W_{m}+ndl_{q}l_{d})$ $O(D_{m}+\log dl_{d})$ $O(W_{m}n)$ $O(D_{m})$ $O(ndl_{d})$
+EMDR (Singh et al., 2021) $O((k+1)W_{m}+nd)$ $O(2D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+Self-RAG (Asai et al., 2024) $O((2k+1)W_{m}+nd+k)$ $O(2D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+Chain-of-Note (Yu et al., 2024b) $O((ks(n)+1)W_{m}+nd)$ $O((ks(n)+1)D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+RAPTOR (Sarthi et al., 2024) $O(W_{m}+nd+s(n)d)$ $O(D_{m}+\log d)$ $O(W_{m}n+W_{m}s(n))$ $O(D_{m})$ $O(nd+s(n)d)$
+RAGraph (Jiang et al., 2024) $O(W_{m}+nd+s(n)d)$ $O(D_{m}+\log d)$ $O(W_{m}+W_{e})$ $O(D_{m}+D_{e})$ $O(nd+s(n)d)$
+RQ-RAG (Chan et al., 2024) $O(s(n)(W_{m}+nd))$ $O(s(n)(D_{m}+\log d))$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+ActiveRAG (Xu et al., 2024) $O(3W_{m}+nd)$ $O(2D_{m}+\log n)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+HiQA (Chen et al., 2024d) $O(W_{m}+nd+s(n)d)$ $O(D_{m}+\log d)$ $O(W_{m}n+s(n)d)$ $O(D_{m})$ $O(nd+s(n)d)$
+GraphRAG (Edge et al., 2025) $O(s(n)(W_{m}+d))$ $O(D_{m}+s(n)d)$ $O(W_{m}n+W_{m}s(n)+W_{e})$ $O(D_{m}+D_{e})$ $O(nd+s(n)d)$
+Fusion RAG (Rackauckas, 2024) $O(s(n)W_{m}+knd)$ $O(2D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+Meta-chunking (Zhao et al., 2025b) $O(W_{m}+(n+s(n))d)$ $O(D_{m}+\log d)$ $O(l_{d}W_{m}+s(n)W_{m})$ $O((l_{d}+1)D_{m})$ $O(nd+s(n)d)$
+MoC (Zhao et al., 2025a) $O(W_{m}+(n+s(n))d)$ $O(D_{m}+\log d)$ $O(W_{m}n+s(n))$ $O(D_{m})$ $O(nd+s(n)d)$
+Parametric RAG (Su et al., 2025) $O(W_{m}+nd+s(n))$ $O(D_{m}+\log d+s(n))$ $O(W_{m}ns(n))$ $O(D_{m}s(n))$ $O(ns(n))$
+SuperRAG (Yang et al., 2025) $O(W_{m}+nd+W_{i})$ $O(D_{m}+\log d+D_{i})$ $O(W_{m}n+s(n)+W_{e})$ $O(D_{m}+D_{e})$ $O(nd+s(n))$
+HiRAG (Huang et al., 2025a) $O(W_{m}+nd+W_{i})$ $O(D_{m}+\log d+D_{i})$ $O(W_{m}n+s(n)+W_{e})$ $O(D_{m}+D_{e})$ $O(nd+s(n))$
+MRAG [This Work] $O(W_{m}+nd)$ $O(D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
+4 Benchmarking Multi-Aspectuality
+To assess how well MRAG performs on multi-aspect queries, we need (1) datasets of multi-aspectual documents, (2) queries to the LLM that require retrieving multi-aspect documents, and (3) metrics that assess how well a RAG scheme retrieves such multi-aspect data.
+We now summarize these three elements; details are in Appendix G.
+We construct three multi-aspect datasets: (1) a synthetic Wikipedia dataset with documents sampled from clearly distinct categories (e.g., countries, shipwrecks, board games); (2) a real-world-based legal document dataset annotated by a legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.); and (3) a real-world-based dataset of industry accident reports, categorized by cause.
+For each dataset, we generate multi-aspect queries using GPT-4o, combining $n$ distinct categories into single queries (with $n\in\{1,2,3,4,5,6,10,15,20,25\}$).
 For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories.
-We create such queries by selecting $n$ categories, sampling a document from each selected category (ensuring there are no duplicates overall), and then generating a story that combines these documents, using an LLM (GPT-3.5 Turbo).
-We construct 25 queries with 1, 5, 10, 15 and 20 aspects (125 queries in total).
-An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 3.
-3.3 Metrics
-We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality.
-For a query $Q$, a used retrieval strategy $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$.
-Then, $S(Q,n)$ is the set of the actually retrieved documents.
-We define the Retrieval Success Ratio as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, i.e., the ratio of successfully retrieved relevant documents.
-Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category.
-While less desired, it still increases chances for a more accurate LLM answer following the retrieval.
-For example, when asking the LLM to determine the cause of an industry accident, fetching the documents in the same category as the accident being queried about, improves the chances for the LLM to give a more relevant answer.
-To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$.
-It has the same form as $\Xi(Q,n)$ above, with one difference: $S(Q,n)$ is now the set of all the retrieved documents that belong to categories of the ideal desired documents.
-Finally, to combine these two metrics, we use the Weighted Retrieval Success Ratio $\Xi_{w}$ as $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$.
-By varying $w$, the user can adjust the importance of exact document matches and category matches.
-An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 3.
-Figure 3: An example query used to evaluate different RAG strategies.
-We mention the documents to be fetched in the text and then assess the success ratio of different RAG strategies in finding these documents and their categories.
-We mark exact document matches Refer to caption , category matches Refer to caption , documents that match a category multiple times Refer to caption , and text segments with no matching document Refer to caption .
-Finally, we show the weighted success ratio for each strategy, taking a 2:1 weighting (prioritizing the exact article matches).
+To evaluate retrieval performance, we introduce three metrics for assessing multi-aspectuality.
+Let $Q$ be a query, $S$ a Reranker scheme, and $Q_{rel}$ the ideal set of $n$ relevant documents.
+The Retrieval Success Ratio is defined as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, measuring the fraction of exactly matched documents.
+The Category Retrieval Success Ratio $\Xi_{c}$ extends this by also accepting matches from the same categories as those in $Q_{rel}$ even if the exact document has not been matched.
+Finally, we define a tunable Weighted Success Ratio $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$, allowing the user to adjust the trade-off between exact and category-level matches via the weight $w$.
+We motivate category-level retrieval using both previous works (V et al., 2025; Liu et al., 2021; Chen et al., 2024a; Xiao et al., 2024) as well as our hands-on experience in legal and industrial accident analysis.
+Namely, even when exact matches are missing, retrieving documents from the same semantic category can enhance generation, as supported by the classic clustering hypothesis in information retrieval (V et al., 2025; Liu et al., 2021): documents in the same “neighborhood” are likely to be relevant to the same query.
+Recent studies further validate this across open-domain QA and ontology-guided RAG (Chen et al., 2024a; Xiao et al., 2024).
 5 Evaluation
-We now illustrate the advantages of MRAG over the state of the art.
-Comparison Baselines We consider three main baselines: Standard RAG, Split RAG, and Fusion RAG (Rackauckas, 2024).
+We now illustrate the advantages of MRAG over the state of the art. Further details on evaluation setup and additional results are in – respectively – Appendix H and I.
+Comparison Baselines.
+We consider three main baselines: Standard RAG, Split RAG, and Fusion RAG (Rackauckas, 2024).
 The first represents a modern RAG pipeline in which each document uses the activations of the last decoder layer as its embedding.
-The second is a blend between Standard RAG and MRAG.
-Specifically, it splits the activation of the last decoder layer in the same way as MRAG and applies a voting strategy.
+The second is a blend between Standard RAG and MRAG; it splits the activation of the last decoder layer in the same way as MRAG and applies a voting strategy.
 The purpose of Split RAG is to show that MRAG’s benefits come from using the multi-head output as embedding and not merely using multiple embedding spaces.
-Additionally, we consider Fusion RAG (Rackauckas, 2024), an optional mechanism that we harness to further enhance the benefits of MRAG at the cost of additional tokens (detailed in Section 4.3).
+Additionally, we consider Fusion RAG (Rackauckas, 2024), an optional mechanism that we harness to further enhance the benefits of MRAG at the tradeoff of additional tokens (detailed in Section 5.3).
+Embeddings & models.
+While MRAG allows for extracting multi-aspect embeddings from any block, we found that the last MHA works best in our experiments.
+MRAG can use any embedding model with MHA, including models such as RetroMAE (Xiao et al., 2022) and the classic BGE-embeddings (Xiao et al., 2022; Chen et al., 2024b).
+We consider two embedding models from the MTEB leaderboard (Huggingface, 2025), the SFR-Embedding-Model (Meng et al., 2024) and the e5-mistral-7b-instruct (Wang et al., 2024), both based on the Mistral 7B architecture with 32 decoder blocks and 32 attention heads.
 We use queries and metrics introduced in Section 4.
 We use the weighted retrieval success ratio with 2:1 weighting, which considers category matches as relevant but prioritizes the exact document matches.
-Figure 3 shows an example query and metrics usage.
-Each query requires retrieving a specific number of documents and the corresponding non-overlapping categories which define the ground truth.
-We fetch the top $k$ documents from a database, where $k$ is the “total number of documents fetched for a tested RAG scheme” (including potentially mismatches).
-Among these $k$ documents, we search for matches with the ground truth.
 Samples & Summaries Each data point in our plots corresponds to 25 queries.
@@
 Our primary focus is on the average retrieval performance among those 25 queries.
-Figure 4: Retrieval success ratio over 25 queries between MRAG and Standard RAG; each query uses 10 different aspects.
-The top part presents exact document matches, the bottom part presents category only matches (we explain the metrics in Section 3).
-A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (the number of documents fetched for each query is 30).
-Figure 5: Relative retrieval improvement of MRAG over Standard RAG across queries with different numbers of aspects and different embedding models (SFR in the left side, e5 in the right side).
-4.1 Superior Performance for Multi-Aspect Queries
-We start from the query example in Figure 3 and show first the absolute retrieval performance of MRAG over Standard RAG in Figure 4.
+5.1 MRAG Delivers Superior Performance for Multi-Aspect Queries
+We start with the Wikipedia multi-aspect dataset and corresponding queries (cf.
+Section 4).
+In each query, we mention the documents to be fetched in the text and then assess the success ratio of different RAG strategies in finding these documents and their categories (a full example of such a query is in Figure 10 in Appendix G).
+We show first the absolute retrieval performance of MRAG over Standard RAG in Figure 3.
 We fix the number of aspects present in the queries to 10, and vary the total number of retrieved documents from 10 to 30.
-MRAG consistently outperforms Standard RAG ($>10\%$ increase in the retrieval success ratio on average for exact document matches).
-Moreover, the retrieval performance increase is even more significant on category matches ($>25\%$ increase in the retrieval success ratio on average).
+MRAG consistently outperforms Standard RAG ($10\%$ increase in the retrieval success ratio on average for exact document matches).
+Moreover, the retrieval performance increase is even more significant on category matches ($25\%$ increase in the retrieval success ratio on average).
 The performance increase is further detailed in the histograms on the right side.
 Here, for a specific number of documents fetched, MRAG’s histogram indicates a better distribution of retrieval success ratios (across all 25 queries).
+This gain stems from MRAG’s ability to decompose query semantics across multiple attention heads, increasing the likelihood of matching each aspect with a semantically aligned document.
+Figure 3: Retrieval success ratio over 25 queries between MRAG and Standard RAG; each query uses 10 different aspects.
+The top part presents exact document matches, the bottom part presents category only matches (we explain the metrics in Sec.
+4).
+A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (there are 30 documents fetched for each query).
 Next, Figure 4 shows the relative weighted performance improvement of MRAG with respect to Standard RAG as we vary the number of aspects present in the queries.
 We show data for two different embedding models (SFR and e5).
-MRAG consistently outperforms the Standard RAG by 10-20% on average, not only across the number of documents fetched, but also across the number of aspects present in the replies, for both models.
-4.2 Comparable Performance for Single-Aspect Queries
-We additionally show in Table 1 that MRAG performs on-par with Standard RAG on queries from our multi-aspect dataset where only a single aspect is expected.
-Hence, our approach does not suffer from significant decrease in performance for single-aspect tasks.
-Table 1: Retrieval success ratio (the exact document match) for 25 queries with a single aspect.
-[…]
-Documents Fetched Multi-Aspect Dataset Legal Dataset Accidents Dataset
+MRAG consistently outperforms the Standard RAG by 10-20% on average, not only across the number of documents fetched, but also across the increasing counts of aspects present in the replies, and does it for both embedding models.
+This robustness suggests that MRAG scales better than Standard RAG with query complexity, as its multi-head representation distributes semantic load more evenly than the single-vector baseline.
+Figure 4: Relative retrieval improvement of MRAG over Standard RAG across queries with different numbers of aspects and different embedding models (SFR in the left side, e5 in the right side).
+To further illustrate advantages of MRAG, we also consider two real-word use cases from in-house industry data analytics projects, namely, the synthesis of legal documents and the analysis of causes of chemical plant accidents.
+The results are in Figure 5.
+In the former (the left side), the task is to create a document based on user requirements that may be related to different aspects, for example to the law being considered (e.g., the British or the US one), the subject (e.g., energetic or civil), the style of the document (e.g., aggressive or mild), and others.
+This task is executed with RAG that can fetch documents from a database.
+In the latter (the right side), the task is to discover a cause of an accident.
+Here, one also wants to retrieve documents from a database that should be used in the LLM context to facilitate discovering the cause of the accident.
+The causes are grouped in categories such as utility impact due to severe weather, lack of preparedness and planning, incorrect installation of equipment, lack of maintenance, et cetera.
+Similarly to the previous analyses, we measure the retrieval success ratio over corresponding databases.
+MRAG offers advantages over other schemes.
+Figure 5: Average improvement of the retrieval success ratio of MRAG and Split RAG over Standard RAG for two real-world workloads constructing legal documents (left) and discovering causes of industry accidents (right).
+Figure 6: Relative retrieval improvements of MRAG over Standard RAG for the SFR embedding model compared with Split RAG (the blue plots), and the relative retrieval improvements of Fusion MRAG over Standard RAG compared with Fusion RAG (the red plots).
+We also delve deeper into the underlying factors for MRAG’s performance gains.
+For this, we compare MRAG to the Split RAG baseline in Figure 6.
+The blue plots show the relative weighted performance of MRAG and Split RAG over Standard RAG.
+MRAG performs better than Split RAG, illustrating that its high accuracy is due to the actual multi-head part, and not merely just partitioning the vector and using multiple embedding spaces.
+5.2 MRAG Ensures High Performance for Single-Aspect Queries As Well
+We additionally show in Table 2 that MRAG performs on-par with Standard RAG on queries where only a single aspect is expected.
+This confirms that MRAG’s design generalizes well: in low-aspect settings, the aggregation over heads still captures dominant semantics, effectively collapsing into a strong single-vector representation with only a negligible accuracy drop for single-aspect tasks.
+Table 2: Retrieval success ratio (the exact document match) for 25 single-aspect queries on different datasets (Multi-Aspect Dataset, Legal Dataset, Accidents Dataset), using different embedding models (SFR, e5).
+For every query, a specific single document (single-aspect) is expected to be among the fetched documents for the retrieval to be classified as successful.
+Documents Fetched Wikipedia Dataset Legal Dataset Accidents Dataset
 SFR e5 SFR SFR
@@
 3 25/25 25/25 25/25 25/25 25/25 25/25 25/25 25/25
-4.3 Further Improvements with Additional Tokens
-We now show that MRAG can be seamlessly integrated with other RAG approaches: We combine MRAG with Fusion RAG, representing RAG schemes that use an LLM (additional token cost) for more accurate retrieval.
+5.3 MRAG Seamlessly Enhances Existing RAG Schemes
+MRAG’s simplicity ensures that it can be seamlessly integrated with other RAG approaches.
+As an example, we combine MRAG with Fusion RAG, which uses an LLM (additional token cost) for even more accurate retrieval.
 Fusion RAG uses an LLM to create a fixed number of questions about the RAG query.
 Each question is separately applied through an embedding model using Standard RAG.
-We apply MRAG’s approach to each of these questions and denote the combined scheme as Fusion MRAG.
-Red plots of Figure 6 show that both Fusion RAG and Fusion MRAG perform better than Standard RAG, on average gaining 10 to 30% in accuracy.
-Fusion MRAG performs consistently better than pure Fusion RAG, indicating that these optimizations can be combined together.
+This relies on multiple LLM calls and heuristic reranking, inflating latency and cost.
+We apply MRAG to each of these questions and denote the combined scheme as Fusion MRAG.
+Red plots of Figure 6 show that Fusion MRAG consistently outperforms pure Fusion RAG, indicating that these optimizations can be combined together.
 However, both Fusion strategies introduce a greater variance than MRAG and additional costs in terms of compute, latency, and tokens.
-4.4 Benefits from Multi-Head Attention Solely
-We also compare MRAG to the Split RAG baseline in Figure 6.
-The blue plots show the relative weighted performance of MRAG and Split RAG over Standard RAG.
-MRAG performs better than Split RAG, illustrating that its high accuracy is due to the actual multi-head part, and not merely just partitioning the vector and using multiple embedding spaces.
-Figure 6: Relative retrieval improvements of MRAG over Standard RAG for the SFR embedding model compared with Split RAG (the blue plots), and the relative retrieval improvements of Fusion MRAG over Standard RAG compared with Fusion RAG (the red plots).
-4.5 Real-World Workloads
-Figure 7: Average improvement of the retrieval success ratio of MRAG and Split RAG over Standard RAG for two real-world workloads constructing legal documents (left) and discovering causes of industry accidents (right).
-To further illustrate advantages of MRAG, we also consider two real-word use cases from in-house industry data analytics projects, namely, the synthesis of legal documents and the analysis of causes of chemical plant accidents.
-The results are in Figure 7.
-In the former (the left side), the task is to create a document based on user requirements that may be related to different aspects, for example to the law being considered (e.g., the British or the US one), the subject (e.g., energetic or civil), the style of the document (e.g., aggressive or mild), etc..
-This task is executed with RAG that can fetch documents from a database.
-In the latter (the right side), the task is to discover a cause of an accident.
-Here, one also wants to retrieve documents from a database that should be used in the LLM context to facilitate discovering the cause of the accident.
-The causes are grouped in categories such as utility impact due to severe weather, lack of preparedness and planning, incorrect installation of equipment, lack of maintenance, etc..
-Similarly to the previous analyses, we measure the retrieval success ratio over corresponding databases.
-MRAG offers advantages over other schemes.
-4.6 Additional Analyses
-Figure 8: Evaluation of different voting strategies for MRAG and Split RAG.
-We also analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one).
-Here, we consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention.
-We discover that the last multi-head attention performs the best when compared with the Standard RAG.
-We also illustrate selected representative data from a long investigation into two additional voting strategies for MRAG.
-We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$.
-Figure 8 shows that these voting strategies perform worse on average than our selected strategy for MRAG, justifying its design and selection (described in Section 2.3).
-We also consider two voting strategies for Split RAG, to further deepen the empirical evaluation.
-Split (1) only uses the exponential lowering of significance ($w_{i,p}=2^{-p}$) and Split (2) which uses the same strategy as MRAG ($w_{i,p}=s_{i}\cdot 2^{-p}$).
-Figure 8 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
-The complexity of the importance score calculation (Algorithm 1) is $\mathcal{O}(n^{2})$ where $n$ is the number of the embedded documents; it is dominated by calculating the pair-wise cosine similarity and the calculation of the norm.
-Please note that this step needs to be done only once for each dataset and is not a bottleneck.
-Finally, in addition to the performance evaluation, we also investigated the attention heads of the SFR-Embedding-Mistral model as well as LLaMA-2 7B model (model not fine-tuned for text-embedding tasks).
-This analysis is presented in Appendix C.
-5 Related Work
-Our work touches on many areas which we now briefly discuss.
-Many RAG schemes appeared recently (Gao et al., 2024), using the output of the last decoder layer for embedding generation.
-In contrast, MRAG leverages different embedding spaces of attention heads to focus on different aspects of documents and queries.
-As such, it can be combined with other schemes to further improve RAG pipelines.
-Retrieval is sometimes enhanced by a cross-encoder reranking phase (Rosa et al., 2022; Nogueira & Cho, 2020; Nogueira et al., 2020; Li et al., 2021; Gao et al., 2021; MacAvaney et al., 2019).
-In such solutions, typically after retrieving a set of relevant chunks, they are re-ranked using specialized models.
-In this work, we focus solely on the first retrieval phase, so MRAG can be seamlessly used in conjunction with such cross-encoders.
-Structure-enhanced RAG schemes employ different strategies for structuring text to improve retrieval quality.
-A common idea is to construct a Knowledge Graph from text, which enables retrieval amongst entities and relationships (Jiang et al., 2024; Delile et al., 2024; Hussien et al., 2024; Bui et al., 2024; Xu et al., 2024a).
-RAPTOR (Sarthi et al., 2024) generates multi-level summaries for clusters of related chunks, building a tree of summaries with increasing levels of abstraction to better capture the meaning of the text.
-Graph RAG (Edge et al., 2024) creates a Knowledge Graph, and summarizes communities in the graph, which provide data at the different levels of abstraction.
-All these systems try to improve RAG quality by utilizing additional structures that describe entity relationships or the inner organization of text.
-Usually, they need a sophisticated preprocessing phase to prepare such structures.
-MRAG achieves the improvement solely based on the embedding model and has no additional storage requirements, and can be combined with any of these schemes.
+5.4 MRAG Seamlessly Enhances Downstream LLM Generation
+We also illustrate that MRAG enhances the downstream LLM generation with its improved multi-aspect retrieval.
+To show this, we use a sampled subset of multi-aspect Wikipedia queries, for which we applied both Standard RAG and MRAG to retrieve supporting documents.
+These documents are then integrated into the prompt templates and are passed to the LLM to answer the original query with the aid of the retrieved data.
+To quantify the effect from RAG, we count facts in the LLM output (e.g., years or named entities such as locations).
+On average, MRAG generations contain on average 15.4 pieces of factual information per query, compared to 11.2 for Standard RAG.
+The average improvement further confirms MRAG’s advantages.
+By explicitly leveraging multi-aspectual embeddings from MHA, one increases the likelihood that the LLM generation includes diverse and complementary facts, especially for complex queries spanning multiple domains (e.g., combining historical events with technological timelines).
+Overall, MRAG helps the LLM in delivering richer, more comprehensive responses.
+5.5 MRAG Ensures No Latency and Storage Overheads
+MRAG introduces no latency overhead at query embedding time, as all head-level embeddings are extracted in parallel from a single standard forward pass.
+Retrieval across embedding subspaces is also fully parallelizable with modern vector databases (Pan et al., 2024; Ma et al., 2025), and our use of a modest number of heads (e.g., 16–32 for SFR-Embedding-Model) ensures this parallelism is within easy reach.
+Moreover, as the total dimensionality of embeddings remains unchanged (e.g., 1024 split across 32 heads), MRAG needs no additional storage compared to standard RAG.
+5.6 Further Analyses & Ablation Studies
+Additional analyses are in Appendix I, they include analyzing the impact of using embeddings from different decoder blocks (rather than the last one), scalability of preprocessing, and additional voting strategies for reranking.
+These analyses all confirm the previous findings of MRAG broadly outperforming other baselines.
+6 Related Work & Advantages of Multi-Head RAG
+RAG Solutions.
+We compare MRAG to a large number of both traditional and modern RAG solutions in Table 3 in terms of design advantages (complexity analysis and empirical evaluation are in, respectively, Sections 3 and 5).
+While prior RAG systems support retrieving multiple documents for a single query (e.g., RAG (Lewis et al., 2020), EMDR (Singh et al., 2021)), none of them generates multi-aspect embeddings per document, and therefore do not offer multi-aspectuality.
+Similarly, although these systems employ Transformer architectures with MHA, they make no use of the MHA internal structure.
+Other methods such as Poly-encoder (Humeau et al., 2020) and ColBERT (Khattab & Zaharia, 2020) do produce multiple embeddings per document, but for a different reason: they are based on models such as BERT, which inherently yields token-level embeddings.
+Thus, these models require multiple vectors per document simply to cover its content at the token level, not to represent distinct semantic aspects.
+In contrast, MRAG leverages the powerful internal structure of modern decoder-based LLMs, where a small number of vectors derived from MHA (or even a single vector from the final decoder block, as in standard RAG) suffices to represent the semantics of an entire document or a chunk, as indicated by recent work (Lee et al., 2025; Besta et al., 2025).
+Building on this foundation, MRAG achieves further practical advantages.
+It introduces scalable preprocessing since (1) global embeddings are computed in a single forward pass per document and (2) scoring of heads is straightforwardly parallelizable.
+At inference time, it incurs no additional cost relative to standard RAG – in contrast to token-level approaches, which require computing numerous pairwise token similarities during retrieval, our method only compares compact, chunk-level vectors.
+Storage overhead is also minimal: the MHA-derived embeddings match the dimensionality of standard last-layer embeddings (detailed time and storage complexity analyses are in Section 3).
+Finally, MRAG requires no training and is highly versatile: it can be applied to any Transformer model with MHA, and is easily extensible to other modalities such as vision and graph data, where Transformer architectures are now common.
+Table 3: Comparison of the advantages of different RAG schemes (sorted top to bottom chronologically).
+No additional training: does a given scheme require additional training beyond standard pre-training and fine-tuning?
+Works with any MHA LLM: does a given scheme can seamlessly work with any multi-head attention (MHA) LLM?
+Extensibility to other modalities: could a given scheme be relatively easily used beyond LLMs, e.g., with Graph Foundation Models (GFMs), Vision Models, and others?
+No overhead at inference: does a given scheme enable zero additional overhead at the inference?
+Scalable preprocessing, low storage overhead: does a given scheme enable scalable preprocessing and little storage overhead, respectively?
+(details in Section 3 and Table 1).
+Multi-Aspectuality: does a given scheme enable extracting multiple aspects of indexed data?
+“”: full support, “”: partial support, “”: no support, “”: unknown.
+Scheme
+No additional
+training
+Works with any
+MHA LLM
+Extensibility to
+other modalities
+No overhead
+at inference
+Scalable
+preprocessing
+Low storage
+overhead
+Multi-
+Aspectuality
+Poly-encoders (Humeau et al., 2020) (BERT based)
+Lewis et al. (2020)
+ColBERT (Khattab & Zaharia, 2020) (BERT based)
+EMDR (Singh et al., 2021)
+Self-RAG (Asai et al., 2024)
+Chain-of-Note (Yu et al., 2024b)
+RAPTOR (Sarthi et al., 2024)
+RAGraph (Jiang et al., 2024) (only GFMs) (only GFMs)
+RQ-RAG (Chan et al., 2024)
+ActiveRAG (Xu et al., 2024)
+HiQA (Chen et al., 2024d)
+GraphRAG (Edge et al., 2025)
+Fusion RAG (Rackauckas, 2024)
+Meta-chunking (Zhao et al., 2025b)
+MoC (Zhao et al., 2025a)
+Parametric RAG (Su et al., 2025)
+SuperRAG (Yang et al., 2025)
+HiRAG (Huang et al., 2025a)
+MRAG [This work]
+Reranking.
+Retrieval is sometimes enhanced by a reranking phase (Rosa et al., 2022; Nogueira & Cho, 2020; Nogueira et al., 2020; Li et al., 2023; Gao et al., 2021; MacAvaney et al., 2019).
+Here, after retrieving a set of relevant chunks, they are re-ranked using specialized models.
+In this work, we provide a heuristic reranker that considers multi-aspectuality, but we design MRAG specifically so that it can be seamlessly used in conjunction with any other existing cross-encoders.
+Multi-Head Embeddings Outside RAG.
+Several methods propose a new model variant or architectural modification to the Transformer in order to better exploit MHA (Park et al., 2020; Huang et al., 2019; Wang et al., 2020; Xue & Aletras, 2023).
+For example, MHSAN (Park et al., 2020) introduces a novel visual-semantic embedding network that extracts multiple region- and phrase-sensitive features using MHA.
+Wang et al.
+(2020) develop a speaker embedding network that explicitly enforces head-level diversity via contrastive learning across resolutions.
+Xue & Aletras (2023) propose PIT, a Transformer variant that composes attention heads across layers to reduce redundancy and enable efficient inference.
+Vashisht et al.
+(2025) introduce MAGE, a technique that mixes heads across models to improve generalization through stochastic head combinations.
+In contrast, MRAG is the first method to harness embeddings from MHA in pre-trained, decoder-based embedding LLMs for the purpose of more effective RAG.
+Unlike the above works, MRAG requires no architectural modifications, no training, and no custom modules, being fully plug-and-play.
 7 Conclusion
-Retrieval Augmented Generation (RAG) is pivotal for democratizing access to accurate and relevant outputs from large language models (LLMs).
-Enhancing the precision and relevance of these outputs is a critical goal, especially given the challenges posed by queries requiring the retrieval of multiple documents with significantly different contents.
-These complex queries are common across various domains, but existing RAG solutions struggle because the embeddings of the necessary documents can be far apart in the embedding space, complicating their retrieval.
-To address this gap, we introduced Multi-Head RAG (MRAG), a novel scheme that leverages the activations from the multi-head attention layer of decoder models instead of the traditional feed-forward layer.
-This approach is grounded in the insight that different attention heads can capture distinct aspects of the data.
+Retrieval Augmented Generation (RAG) is crucial for democratizing access to accurate and relevant outputs from large language models (LLMs).
+However, enhancing the precision of RAG is non-trivial, especially given the challenges posed by queries requiring the retrieval of multiple documents with significantly different contents.
+These complex multi-aspectual queries are common across various domains, but existing RAG solutions struggle with them because the embeddings of the necessary documents can be far apart in the embedding space, complicating their retrieval.
+To address this issue, we introduce Multi-Head RAG (MRAG), a novel RAG scheme based on a simple yet powerful idea: leverage the activations from the multi-head attention (MHA) layer of decoder models instead of the traditional feed-forward layer.
+This approach is based on the insight, confirmed by an extensive literature survey, that different attention heads can capture distinct aspects of the data even without additional training.
 By using these diverse activations, MRAG creates embeddings that better represent the multifaceted nature of data items and queries, thus enhancing the retrieval accuracy for complex, multi-aspect queries.
-The simplicity and versatility of this idea allow it to be seamlessly integrated into any modern RAG pipeline or data analytics framework.
-Our comprehensive evaluation methodology, including specific metrics, synthetic datasets, and real-world use cases, demonstrates MRAG’s effectiveness.
+The simplicity and versatility of this idea allow it to be seamlessly integrated into any modern RAG pipeline or data analytics framework, and to be applied to other classes of models beyond LLMs.
+Our comprehensive evaluation methodology for multi-aspect queries, including novel metrics, synthetic datasets, and real-world use cases, demonstrates MRAG’s effectiveness.
 The results indicate a significant improvement in the relevance of retrieved documents, with up to 20% better performance compared to modern RAG baselines.
+Moreover, harnessing MRAG results in enhanced downstream LLM generation by providing richer and more factual LLM outputs.
 This validates MRAG’s potential to handle the intricacies of multi-aspect queries effectively.
-Moreover, MRAG proves to be both cost-effective and energy-efficient.
-It does not require additional LLM queries, multiple model instances, increased storage, or multiple inference passes over the embedding model.
-This efficiency, combined with the enhanced retrieval accuracy, positions MRAG as a valuable advancement in the field of LLMs and RAG systems.
-By addressing the challenges of multi-aspectuality in queries, MRAG paves the way for more reliable and accurate LLM applications across diverse industries.
-Acknowledgements
-We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines, and for their excellent technical support.
+Our analysis of nearly 20 RAG baselines illustrates that MRAG is both highly effective and efficient because – unlike other RAG schemes – it does not require additional LLM queries, multiple model instances, increased storage, additional training, or multiple inference passes over the embedding model.
+These advantages, combined with the enhanced retrieval accuracy and MRAG’s scalability, make MRAG a valuable design in the field of LLMs and RAG systems.
+Acknowledgments
+We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault, Daint and Alps machines, and for their excellent technical support.
 We thank Timo Schneider for help with infrastructure at SPCL.
@@
 PLG/2024/017103.
+We thank Kurt Krieg for useful insights into the nature of processing plant accidents.
 Appendix
-Appendix A Model Design: Additional Details
-A.1 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
+Appendix A Review of Multi-Aspectuality in Industry
+Real-world decision-making and analysis tasks often require a holistic integration of multiple, semantically distinct information sources.
+Literature from diverse domains—from safety investigations to diagnostics and enterprise analytics—consistently emphasizes that no single data stream suffices for complex tasks.
+Instead, success comes from combining heterogeneous elements (design details, human factors, environmental context, etc.) into a unified understanding.
+Below we highlight several examples that support this multi-aspectual reasoning framework.
+In safety-critical environments, accident and incident investigations rely on aggregating information from disparate sources.
+For example, a U.S.
+Federal Railroad Administration report stresses that collecting multiple sources of information is essential for effective event reconstruction (Office of Railroad Safety, 2014).
+The Columbia Accident Investigation Board similarly combined telemetry, debris forensics, environmental data, and even autopsy reports to reconstruct the chain of events behind the Space Shuttle Columbia disaster (Columbia Accident Investigation Board, 2003).
+These investigations exemplify the need to integrate technical, procedural, and human-centered aspects of data to obtain actionable conclusions.
+In medicine, particularly oncology, the concept of integrated diagnostics exemplifies multi-aspect decision-making.
+Physicians routinely combine patient histories, radiological scans, lab values, pathology slides, and genetic tests to form a diagnosis.
+This is no longer optional: as modern datasets become more complex, integrating semantically distinct modalities has become necessary to reach accurate, personalized outcomes (Messiou et al., 2023; Kalia, 2013).
+The same applies to enterprise settings.
+For example, airport operations management integrates foot traffic sensors, weather feeds, and gate schedules – among others – to optimize personnel allocation and prevent congestion (Zografos et al., 2013).
+Similarly, in legal and financial firms, cross-silo systems integrate internal metrics (e.g., billing and staffing) with external sources (e.g., news, contracts, social media) to guide decision-making and strategic planning (OpenText, 2022).
+These examples show that modern organizational workflows demand the integration of multiple semantically distinct data sources.
+In summary, across domains like industrial safety, healthcare, and business intelligence, it is widely recognized that multi-aspectuality—combining diverse, independently relevant information fragments—is essential for accurate and effective decision-making (Office of Railroad Safety, 2014; Columbia Accident Investigation Board, 2003; Reason et al., 2006; Kalia, 2013; Packham, 2017; Messiou et al., 2023; Multer et al., 2013; Coury et al., 2010; Harle, 1997; O’Hare, 2000; Gordon et al., 2005; Bridger, 2021; OpenText, 2022).
+In all these cases, the embeddings of documents from such divergent subdomains would be far away from one another in the embedding space when using standard RAG pipelines (as also confirmed by our own datasets in legal and plant accident use cases, see Section 5), underlying the relevance of MRAG.
+Appendix B Mathematical Formulation: Additional Details
+We omit, for clarity, unnecessary details such as layer normalizations. The output of attention head $h$ for the $i$th token $x_{i}$ is defined as follows (Vaswani et al., 2017):
+$\text{head}^{h}(\mathbf{x}_{i})=\sum_{j}w_{ij}\mathbf{v}_{j}^{h},\quad\text{where}$ (1)
+$\vskip-5.0ptw_{ij}=\text{softmax}\left(\left(\mathbf{q}_{i}^{h}\right)^{T}\mathbf{k}_{j}^{h}\right),\quad\mathbf{q}_{i}^{h}=\mathbf{W}_{q}^{h}\mathbf{x}_{i},\quad\mathbf{k}_{j}^{h}=\mathbf{W}_{k}^{h}\mathbf{x}_{j},\quad\mathbf{v}_{j}^{h}=\mathbf{W}_{v}^{h}\mathbf{x}_{j}$ (2)
+where $\mathbf{W}_{q}^{h},\mathbf{W}_{k}^{h},\mathbf{W}_{v}^{h}$ are, respectively, learnable query, key, and value projections associated with head $h$, and $\mathbf{x}_{j}$ is the vector embedding of the $j$th token $x_{j}$.
+These outputs get combined to form the output of the $i$th multi-head attention block as follows:
+$\text{MHA}(\mathbf{x}_{i})=\mathbf{W}_{o}\ \text{concat}(\text{head}^{1}(\mathbf{x}_{i}),\dots,\text{head}^{h}(\mathbf{x}_{i}))^{T}$ (3)
+where matrix $\mathbf{W}_{o}$ is the linear layer that combines the outcomes of all attention heads.
+B.1 Standard RAG Embedding
+In standard RAG, a single embedding vector $\mathbf{e}_{\text{std}}\in\mathbb{R}^{d}$ is computed for a chunk by extracting the decoder output for the final token $x_{n}$ after the final feed-forward layer:
+$\mathbf{e}_{\text{std}}=\text{FFN}\left(\text{MHA}(x_{n})\right)$ (4)
+B.2 Multi-Head RAG Embedding
+Instead of compressing all head outputs into a single embedding, MRAG leverages the individual output of each head on the final token $x_{n}$:
+$\mathcal{S}=\{\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})\in\mathbb{R}^{d/h}\mid k=1,\dots,h\}$ (5)
+This results in $h$ head-wise embeddings per chunk, capturing diverse semantic aspects.
+Crucially, this design avoids any increase in memory or compute during inference, as head-level vectors are computed as part of the standard MHA process.
+Appendix C System Design & Implementation: Additional Details
+We provide addition details on system design and implementation.
+C.1 Ranking Strategy Details
+The scoring of embedding spaces is detailed in Algorithm 1.
+Algorithm 1 Importance scores for heads.
+for each head $h_{i}$ do
+  $a_{i},b_{i},count\_a_{i},count\_b_{i}\leftarrow 0$
+  for each embedding $e_{ij}$ in $h_{i}$ do
+[…]
+   $count\_a_{i}\leftarrow count\_a_{i}+1$
+   for subset of $m$ embeddings $e_{ih}$ sampled uniformly at random do
+    $b_{i}\leftarrow b_{i}+\text{cosine-distance}(e_{ij},e_{ih})$
+    $count\_b_{i}\leftarrow count\_b_{i}+1$
+   end for
+  end for
+  $a_{i}\leftarrow a_{i}/count\_a_{i}$; $b_{i}\leftarrow b_{i}/count\_b_{i}$
+  $s_{i}\leftarrow a_{i}\cdot b_{i}$
+end for
+C.2 Ranking Strategy Details
+The voting strategy used by MRAG in its reranker is pictured in Algorithm 2.
 Algorithm 2 Voting strategy.
@@
 for each head $h_{i}$ and its score $s_{i}$ do
-     find best matching $k$ text chunks
-     for each chunk $d_{i,p}$ with index $p$ in top $k$ do
+  find best matching $c$ text chunks
+  for each chunk $d_{i,p}$ with index $p$ in top $c$ do
    $w_{i,p}\leftarrow s_{i}\cdot 2^{-p}$
@@
 end for
-  sort $l$ using weights $w_{i,p}$; return top $k$ elems
-Appendix B Evaluation Methodology: Additional Details
-B.1 Compute Resources
-Our experiments were executed with compute nodes containing 4x NVIDIA GH200 and a total memory of 800 GB.
-In general one GPU with at least 40GB of memory should suffice.
-We used at most 50GB of storage and the OpenAI API as an external resource.
-The full experiments took at most three hours of GPU time and the cost for the OpenAI API were at most \$15.
-We carried out additional experiments, which amounted to around 20 hours of GPU time and cost of \$25 for the OpenAI API.
-Additional evaluation was executed with a mix of compute resources including NVIDIA A100 and V100 GPUs.
-B.2 Dataset Details
-Table 2: Overview of the structure and the number of documents in the respective datasets.
-dataset #categories #topics #documents total #documents
-Wikipedia 25 50 documents per category 1250
-Legal Documents 25 25 per category 10 per topic 6250
-Accident Reports 25 25 per category 10 per topic 6250
-B.3 Prompt Template for the Synthetic Dataset Generation
-Table 3: Prompt template for query generation.
-Please create a story about the attached <number of articles> articles on the topics <list of titles>.
+ sort $l$ using weights $w_{i,p}$
+ return top $k$ elements of $l$
+C.3 Integration with Data Stores
+MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches.
+It can be combined with both the exact and the approximate NN to find the matching (embedding, chunk)-pairs.
+These two parts of the broader RAG processing pipeline are orthogonal to MRAG.
+Appendix D Specification of Prompts
+D.1 Prompt Template for Reader
+Prompt Template for Reader You are presented with a series of articles, each potentially addressing different aspects of a topic.
+1. <article title>
+[<metadata>]
+<body>
+2. <article title>
+[<metadata>]
+<body>
+Task: Carefully analyze the articles.
+When formulating your response to the question below, identify the relevant aspects or claims made within each article.
+Construct your answer by comparing, contrasting, or synthesizing these points in a coherent and logically structured manner.
+Your response should be supported by specific references to the content of the articles.
+Where applicable, acknowledge differences in perspectives, data, or assumptions across the sources.
+Aim for clarity, precision, and concise reasoning grounded in the evidence provided.
+Please answer the following question: <query text>
+D.2 Prompt Template for the Synthetic Dataset Generation
+Prompt Template for Query Generation Please create a story about the attached <number of articles> articles on the topics <list of titles>.
 It is very important that each of the attached articles is relevant to the story, in a way that references the content of the article, not just its title.
@@
 Again, make sure that you reference all the following topics in your story: <list of titles>
-Appendix C Attention Head Analysis
+Appendix E Multi-Aspectuality with Attention Heads without Additional Training
+In MRAG, we extract embeddings from the hidden representations immediately after the attention block in the last decoder layer, avoiding any fine-tuning.
+This decision is based on an existing hypothesis that attention heads in Transformer models naturally differentiate during training, each attending to distinct aspects of the input data distribution.
+E.1 Literature Survey
+This hypothesis has been substantiated across various Transformer families.
+Wang et al.
+(2025a) introduce the Local Learning Coefficient (LLC), a measure of training dynamics at the head level.
+They show that attention heads begin with similar behavior but quickly diverge into functionally distinct clusters, each specializing in different patterns, ranging from local structure to multigram token groups.
+Olsson et al.
+(2022) identified “induction heads” in GPT-style models: specific heads that attend to earlier repeated sequences (e.g., in patterns like “X … Y … X”), enabling the model to learn simple in-context repetition without additional supervision.
+Further studies observed heads that consistently attend to names, suppress repetition, or shift attention predictably to structurally aligned tokens (McDougall et al., 2024; Gould et al., 2024).
+Research in BERT-style models further supports these trends.
+Clark et al.
+(2019) demonstrate that different heads attend to direct objects, nominal modifiers, or punctuation tokens.
+Kovaleva et al.
+(2019) identify broad attention patterns like “vertical” heads (focusing on special tokens) and “diagonal” heads (attending to adjacent words).
+Htut et al.
+(2019) show that many heads correlate with syntactic dependency arcs.
+Notably, while many heads specialize in meaningful ways, others appear to contribute little to model performance.
+Voita et al.
+(2019) and Michel et al.
+(2019) show that a large fraction of heads can be pruned without substantial performance loss, suggesting that specialization tends to concentrate in a smaller subset of effective heads.
+E.2 Analysis of Multi-Head Patterns
 We investigated the attention heads of two models in detail: LLaMA-2 7B and SFR-Embedding-Mistral.
@@
 Still, we encountered two distinct patterns.
-First, the diagonal lines in Figures 9(a) and 9(b) indicate that, when processing a certain input token x, elevated attention is paid to some tokens that came a constant numbers of steps before x.
+First, the diagonal lines in Figures 7(a) and 7(b) indicate that, when processing a certain input token x, elevated attention is paid to some tokens that came a constant numbers of steps before x.
 We postulate that this pattern is likely beneficial to understanding the overall rhythm of a natural language, allowing the model to better identify which words are semantically connected, and which parts of the input text refer to each other.
-Second, horizontal and vertical lines in Figures 9(c) and 9(d) show that these heads learned to pay attention to specific tokens, regardless of how far apart they are within the input sequence.
+Second, horizontal and vertical lines in Figures 7(c) and 7(d) show that these heads learned to pay attention to specific tokens, regardless of how far apart they are within the input sequence.
 An intuitive justification for such patterns is the focus on certain semantic aspects of the input sequence.
@@
 Figure 8: Attention scores for selected attention heads of the LLaMA-2 7B and SFR-Embedding-Mistral models.
-We also detail attention scores (after applying softmax) of selected heads in Figures 10 and 11, when the model is processing the last token of its input.
+We also detail attention scores (after applying softmax) of selected heads in Figures 8 and 9, when the model is processing the last token of its input.
 We see that some tokens gather a lot of attention from most heads, yet there is always a plethora of passages which are attended differently by any two attention heads.
@@
 Figure 9: Attention scores for all attention heads of the SFR-Embedding-Mistral model.
+Appendix F Complexity Analysis: Additional Details
+We now discuss additional details for our complexity analysis of the RAG schemes.
+In Poly-encoders (Humeau et al., 2020) retrieval, we use $s(n)$ to represent additional attention evaluations needed to prepare the $m$ code embeddings and their final context embeddings with all $n$ documents.
+In Lewis et al. (2020) retrieval, the work and depth are increased by the additional generator model evaluation.
+In ColBERT (Khattab & Zaharia, 2020) retrieval, the work is increased linearly and depth logarithmically with the average query and document size due to the maxsim evaluations.
+For storage, the overhead scales linearly with the average size of the document, due to per-token embeddings.
+In EMDR (Singh et al., 2021) retrieval, the work is increased $k+1$ times due to T5 encoders run on the top $k$ documents.
+This corresponds to a $D_{m}$ depth increase as the evaluations can be parallelized.
+In Self-RAG (Asai et al., 2024), once the model decides to conduct retrieval using the retrieval token, it embeds the query, runs nearest $k$ search, and evaluates the model twice for each best-$k$ document to predict the issup, isrel, isuse labels.
+The depth increases only by $D_{m}$ as the evaluations can be parallelized.
+In Chain-of-Note (Yu et al., 2024b), during retrieval the $k$ best documents are fetched and then each is summarized sequentially by generating notes, resulting in a $ks(n)$ work and depth increase, with $s(n)$ representing the cost to generate an average note.
+In RAPTOR (Sarthi et al., 2024), preprocessing involves creating a document tree, with leaves representing documents, and parents containing summaries of children, a cost we represent with $s(n)$ both for work and storage.
+Retrieval involves traversing the tree, which we encapsulate in a different $s(n)$.
+RAGraph (Jiang et al., 2024) preprocessing involves embedding the documents by a graph model and creating the toy graphs, which we account for using $W_{e}$ and $D_{e}$.
+The toy graphs also create additional space requirements, which we account for using $s(n)$.
+For retrieval, additional key information such as environment or position-aware codes is used, represented by another $s(n)$ in our notation.
+In RQ-RAG (Chan et al., 2024), the retrieval may be decomposed into multiple separate queries that are generated by the model and then evaluated in sequence using standard techniques, resulting in increases in work and depth that we denote using $s(n)$.
+In ActiveRAG (Xu et al., 2024), retrieval is conducted using standard techniques, followed by three separate model evaluations: the Knowledge Assimilation (KA), Self-Inquiry (Q), and Thought Accommodation (TA) agents.
+As TA is similar to the work of an LLM answering the query, we omit its cost, like in all other frameworks, resulting in triple the work.
+Q and KA can be evaluated in parallel, increasing the depth only twice.
+HiQA (Chen et al., 2024d) combines multiple retrieval strategies, using vector similarity matching, elastic search with BM25, and keyword matching.
+We represent these with $s(n)$ for work.
+Note that as these are independent, the depth is not increased.
+For preprocessing, HiQA uses a Hierarchical Contextual Augmentor, which creates a data hierarchy and introduces an overhead we denote by $s(n)$.
+HiQA also stores more information alongside vectors, such as keywords, increasing requirements by $s(n)$.
+GraphRAG (Edge et al., 2025) creates a knowledge graph during preprocessing, which we denote with $W_{e}$ and $D_{e}$.
+It then extracts communities and their summaries, which we account for using $s(n)$.
+These are stored and require additional space also denoted by a different $s(n)$.
+During retrieval, GraphRAG uses an LLM to rank all communities in parallel by how useful they are which we estimate using $s(n)$.
+In Fusion RAG (Rackauckas, 2024) retrieval, $k$ queries are generated and for each a standard RAG is evaluated, together with a reranking achieved by another model evaluation, which we estimate as $s(n)$.
+In Meta-chunking (Zhao et al., 2025b), the documents are preprocessed by splitting them into chunks based on perplexity or margin sampling.
+As the decision whether to split or combine sentences in a document is based on an LLM, the preprocessing requires $l_{d}$ evaluations with some $s(n)$ postprocessing.
+The storage requirements are also increased by a different $s(n)$ as documents are stored not as a single vector but multiple vectors based on chunks.
+For the same reason, retrieval requires more work, as the number of vectors is increased by $s(n)$.
+In MoC (Zhao et al., 2025a), chunks are created based on different granularities.
+Similarly to Meta-chunking, this means a larger number of vectors resulting in increased storage requirements, and retrieval work by $s(n)$.
+As MoC also includes routing and meta-chunkers, the preprocessing cost is increased by $s(n)$.
+In Parametric RAG (Su et al., 2025), documents are represented as deltas of parameters that can be applied to the model.
+During preprocessing, the model is fine-tuned on a given document, resulting in a considerable increase of $s(n)$ in work and depth.
+As parameters are considerably larger than the hidden dimension, storage is also increased by a different $s(n)$.
+During retrieval, standard RAG fetches the documents, and the original model needs to be updated with their parameters, increasing work and depth by $s(n)$.
+SuperRAG (Yang et al., 2025) first embeds the documents in a knowledge graph, which we represent by $W_{e}$ and $D_{e}$, and then uses this knowledge graph in retrieval to index it using $W_{i}$ and $D_{i}$.
+These also include any reranking SuperRAG might do.
+HiRAG (Huang et al., 2025a) creates a hierarchical knowledge graph used for indexing in retrieval.
+Similarly, to SuperRAG, this increases the preprocessing and retrieval costs.
+We include in these the additional description and report generation that HiRAG conducts.
+Appendix G Full Specification of Benchmarking Multi-Aspectuality
+We provide more details on how to benchmark multi-aspectuality in RAG.
+Figure 10 shows an example query and metrics usage.
+Each query requires retrieving a specific number of documents and the corresponding non-overlapping categories which define the ground truth.
+We fetch the top $k$ documents from a database, where $k$ is the “total number of documents fetched for a tested RAG scheme” (including potentially mismatches).
+Among these $k$ documents, we search for matches with the ground truth.
+Figure 10: An example query used to evaluate different RAG strategies.
+We mention the documents to be fetched in the text and then assess the success ratio of different RAG strategies in finding these documents and their categories.
+We mark exact document matches Refer to caption , category matches Refer to caption , documents that match a category multiple times Refer to caption , and text segments with no matching document Refer to caption .
+Finally, we show the weighted success ratio for each strategy, taking a 2:1 weighting (prioritizing the exact article matches).
+G.1 Multi-Aspect Datasets
+We first select conceptually different categories of documents for a synthetic dataset.
+Here, we harness publicly available Wikipedia articles.
+In the dataset construction pipeline, the user selects a given number of categories (e.g., countries, board games, historical swords, shipwrecks, etc.) and then, for each category, they sample a specified number of documents.
+The first part of the document (overview) is used as a text chunk to be embedded.
+We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes.
+We also use multi-aspect real-world inspired datasets consisting of NDAs and reports describing industry accidents in chemical processing plants.
+We ensure the usefulness of these datasets by working directly with tech leaders from 3 corporations that rely on RAG in their in-house LLM-driven report generation and analytics frameworks.
+Example categories of the legal documents are legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.).
+Examples of accident causes are natural disasters, human mistakes, or lack of proper training.
+We fully release these datasets to propel RAG research.
+Details on all three datasets can be found in the Appendix H.2.
+In our evaluation, we use a total of 16,500 documents.
+G.2 Multi-Aspect Query Generation
+We also require queries that touch upon a given number of $n$ aspects.
+For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories.
+We create such queries by selecting $n$ categories, sampling a document from each selected category (ensuring there are no duplicates overall), and then generating a story that combines these documents, using an LLM (GPT-4o).
+We construct 160 queries with 1, 2, 3, 4, 5, 6, 10, 15, 20 and 25 aspects (1600 queries in total).
+An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 10.
+G.3 Metrics
+We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality.
+For a query $Q$, a used Reranker scheme $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$.
+Then, $S(Q,n)$ is the set of the actually retrieved documents.
+We define the Retrieval Success Ratio as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, i.e., the ratio of successfully retrieved relevant documents.
+Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category.
+While less desired, it still increases chances for a more accurate LLM answer following the retrieval.
+For example, when asking the LLM to determine the cause of an industry accident, fetching the documents in the same category as the accident being queried about, improves the chances for the LLM to give a more relevant answer.
+To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$.
+It has the same form as $\Xi(Q,n)$ above, with one difference: $S(Q,n)$ is now the set of all the retrieved documents that belong to categories of the ideal desired documents.
+Finally, to combine these two metrics, we use the Weighted Retrieval Success Ratio $\Xi_{w}$ as $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$.
+An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 10.
+Appendix H Evaluation Setup: Additional Details
+H.1 Compute Resources
+Our experiments were executed with compute nodes containing 4x NVIDIA GH200 and a total memory of 800 GB.
+In general one GPU with at least 40GB of memory should suffice.
+We used at most 50GB of storage and the OpenAI API as an external resource.
+The full experiments took at most three hours of GPU time and the cost for the OpenAI API were at most \$15.
+We carried out additional experiments, which amounted to around 20 hours of GPU time and cost of \$25 for the OpenAI API.
+Additional evaluation was executed with a mix of compute resources including NVIDIA A100 and V100 GPUs.
+H.2 Dataset Details
+Table 4: Overview of the structure and the number of documents in the respective datasets.
+dataset #categories #topics #documents total #documents
+Wikipedia 80 50 documents per category 4000
+Legal Documents 25 25 per category 10 per topic 6250
+Accident Reports 25 25 per category 10 per topic 6250
+Appendix I Evaluation: Additional Results
+We provide additional empirical evaluation results.
+I.1 Harnessing Different Decoder Blocks
+We analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one).
+Here, we consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention.
+We discover that the last multi-head attention performs the best when compared with the Standard RAG.
+I.2 Analyzing Different Voting Strategies
+We also illustrate selected representative data from a long investigation into two additional voting strategies for MRAG.
+We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$.
+Figure 11 shows that these voting strategies perform worse on average than our selected strategy for MRAG, justifying its design and selection (described in Section 2.3).
+We also consider two voting strategies for Split RAG, to further deepen the empirical evaluation.
+Split (1) only uses the exponential lowering of significance ($w_{i,p}=2^{-p}$) and Split (2) which uses the same strategy as MRAG ($w_{i,p}=s_{i}\cdot 2^{-p}$).
+Figure 11 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
+Figure 11: Evaluation of different voting strategies for MRAG and Split RAG.
+I.3 Analyzing Preprocessing Overhead
+One-time head importance scoring in MRAG introduces minimal preprocessing overhead on top of the standard embedding scheme.
+The scoring consists of computing: (i) average L2 norms per embedding space, and (ii) average pairwise cosine distances among embeddings within each space.
+To assess practical overhead, we analyze six datasets using the SFR-Embedding-Model.
+The additional time required to compute importance scores is measured as a percentage of the original embedding time.
+For example, on SciFact, the overhead was just 2.7%, and on NFCorpus, only 1.75%.
+Even for moderate-scale corpora such as ArguAna (310 seconds total encoding time), the overhead remained under 5%.
+These results assume full pairwise distance computation across all chunks; in practice, the harnessed sampling makes them even lower.